### v1

# Data Quality Issues in Flare Prediction using Machine Learning Models

Abstract
Machine learning models for forecasting solar flares have been trained and tested using a variety of data sources, such as Space Weather Prediction Center (SWPC) operational and science-quality data. Typically, data from these sources is minimally processed before being used to train and validate a forecasting model. However, predictive performance can be impaired if defects in and inconsistencies between these data sources are ignored. For a number of commonly used data sources, together with softwares that query and then output processed data, we identify their respective defects and inconsistencies, quantify their extent, and show how they can affect the predictions produced by data-driven machine learning forecasting models. We also outline procedures for fixing these issues or at least mitigating their impacts. Finally, based on our thorough comparisons of the impacts of data sources on the trained forecasting model in terms of predictive skill scores, we offer recommendations for the use of different data products in operational forecasting.
Keywords: solar flares, data quality, machine learning, operational forecasting.
1 Introduction
In recent years, machine learning methods have been applied with great success to a variety of space weather forecasting problems, such as solar eruption forecasting (Leka:2018b; leka2019comparison; Camporeale:2019a; chen2019iden; kasapis2022interpretable; whitman2023review; chen2025decent), geomagnetic index prediction (iong2022new; iong2024sparse; chen2025geodgp), and the prediction of terrestrial impacts of space weather events (wang2023forecast; sun2025matrix). These methods have also been used for solar flare forecasting, for which they have improved the state of the art (jiao2020solar; sun2021improved; sun2022predicting; sun2023tensor; chen2024solar). A machine learning method predicts the value of a response variable from the values of one or more predictor variables. In solar flare forecasting in particular, the response is typically the flare class or maximum soft X-ray flux during a flare; the predictors may include remove sensing images such as those from the Solar Dynamics Observatory (SDO) (Boerner:2012a; lemen2012atmospheric; pesnell2012; woods2012; sun2023tensor) or extracted features like the Space eather HMI Active Region Patch (SHARP) parameters (SHARPs). The method is trained to generate predictions via minimization of a loss function that measures the degree to which predicted outcomes deviate from observed outcomes; during training, the mapping from predictors to predictions is manipulated (e.g., by gradient descent) to decrease the loss. Common choices of loss function in the flare forecasting literature include the cross-entropy loss (for binary classification of flares as strong or weak, e.g., chen2019iden), sum of squares loss (for prediction of the logarithm of the peak flux, e.g., jiao2020solar), and the negative log-likelihood (in either classification or regression tasks with statistical models, e.g., sun2023tensor). State-of-the-art binary classification methods for forecasting flares include discriminant analysis (leka2018nwra), logistic regression, support vector machines (bobra2015sola), random forests, gradient boosting machines, long-short term memory (LSTM) neural networks (liu2019predicting), and other more advanced network structures. The literature shows that the performance of machine learning methods is more constrained by the train/validation/test design (random split versus chronological split, for example) and sample size (collecting non-overlapping flare events versus rolling window definition of maximum flaring events) than by the particular machine learning method adopted, due to solar flare data sparsity; see Wang2020lstm; sun2022predicting; chen2024solar for more discussion. Therefore, in this paper, instead of exhaustively evaluating all available machine learning methods, we select representative ones (to be detailed later) and focus on using them to calibrate the impact of the data sources (from different databases) and quality (scientific or operational) we use for training the machine learning model.
Predictors ($X\in\mathbb{R}^{h\times w\times c}$)Refer to caption
$\dots,$
ABSNJZHTOTUSJH
ABSNJZHTOTUSJH
$\dots,$
Predictors ($X\in\mathbb{R}^{d}$)
Method
Prediction ($\widehat{Y}\in\{\text{Flare},\text{No Flare}\}$)
[…]
$\widehat{Y}\in\mathbb{R}$
Prediction ($\widehat{Y}\in\mathbb{R}$)
Figure 1: A schematic illustrating how a machine learning method produces predictions from predictors. Predictors are commonly computed from either images of active regions or summary statistical parameters on those regions. The upper left part of the schematic depicts predictors computed from HMI and AIA images, which can be represented as tensors with dimensions $h\times w\times c$, where $h$, $w$, and $c$ represent the height, width, and channel counts, respectively. For both kinds of images, $h,w=4096$, with $c=1$ for HMI images and $c=10$ for AIA images. The lower left part depicts predictors computed from SHARP parameters, which are summary statistics calculated from HMI images. If the method performs classification, the outcome is the indicator of whether a flare will occur; if the method performs regression, the outcome is the future peak soft X-ray flux. The upper right and lower right show comparisons of actual values to mock predictions for classification and regression, respectively. In the upper right, the green curve represents mock flare probabilities; a classifier may output predicted classes or predicted class probabilities. The images and plots display data from around the time of an X8.7-class flare that occurred on 14 May 2024 in NOAA active region 13664 (HARP 11149).
Different pairs of predictor and outcome types have been used in solar flare forecasting literature. In Figure 1, we provide an overview of how predictions are produced by a machine learning method from different kinds of predictor data for both classification and regression in solar flare forecasting. Predictors are commonly computed from either images of active regions or summary statistics on those regions. The upper left part of the schematic depicts predictors being computed from Helioseismic and Magnetic Imager (HMI) and (Atmospheric Imaging Assembly) AIA images, while the lower left part depicts predictors being computed from SHARP parameters, which are summary statistics calculated from HMI images. If the method performs classification, then the outcome is the binary variable of whether a flare will occur; if the method performs regression, then the outcome is a number, like the future peak soft X-ray flux. The upper and lower right panels show comparisons of actual outcomes to mock predictions for classification and regression, respectively. The images and plots display data from around the time of an X8.7-class flare that occurred on 14 May 2024 in NOAA active region 13664 (HARP 11149). Here we briefly discuss some relevant works corresponding to the parts illustrated in Figure 1. Detailed numerical results will be presented in Sections 2 and 4.
Image Predictors, Binary Response
Image predictors typically consist of multi-channel AIA and HMI images of the Sun. Full-disk images of the Sun at their native resolution of $4096\times 4096$ for each channel are usually too computationally and memory-intensive to be used as inputs to machine learning models, even when performing batch learning. However, resizing these images through down-sampling or other interpolation methods can degrade spatial perception and erase meaningful, fine-grained details that are useful for flare prediction.
On the other hand, multi-channel images focused on active regions [Jin et al., 2024, in prep] must not only deal with potential concurrent temporal effects (for active regions whose timelines overlap) but also with varying image resolutions, as the sizes and shapes of active regions differ. To address this, works such as Pandey2024 perform a rectangular crop that maximizes unsigned flux (a SHARP parameter). However, other work, such as sun2023tensor, uses a similar but different approach by performing a crop centered on the polarity inversion line (PIL). PILs are typically the main source of flare activity, as shown in many works e.g. Zirin:1993; Green:2018; sun2021improved, so such crops can not only keep predictive power high but also allow for succinct comparisons of features (such as topological and spatial features) the model detects with the derived PIL for potential interpretability results.
Using these multi-channel images of active regions is quite common for flare prediction. In the most basic case, we can develop and train machine learning models to predict whether or not a flare will be “strong” within some time horizon. In this case, the response is binary, with “strong” typically meaning C-class or higher (C+) or M-class or higher (M+). While this kind of forecasting may seem simpler than the kinds described below, it can still be difficult. For example, if only M+ flares are considered strong, forecasting can be difficult because of class imbalance and because of the lack of a clear distinction between flares just below and just above the M-class boundary. Instead of just predicting whether a flare will be strong, we can predict the flare class itself. However, this exacerbates the boundary issue, as it now occurs at multiple boundaries.
Many works use image predictors to predict binary flare strength or flare classes. We briefly review a subset of these. Focusing on works that use only 2D magnetogram imaging data, Park2018 trains a deep Convolutional Neural Network (CNN) on full-disk SDO/HMI magnetograms for binary predictions of C+ flares. Meanwhile, huang2018deep and zheng2019solar train CNNs to do multi-class prediction of C-, M-, and X-class flares using full-disk magnetograms. On the other hand, li2022knowledge and sun2022predicting train CNNs on resized HMI magnetograms focused on active regions. Beyond using only 2D magnetogram data, jonas2018flare uses Gabor filters and other techniques on a subset of the AIA imaging channels to train a linear model to predict M+ binary flare outcomes. Additionally, nishizuka2018deep trains a deep neural network on AIA imaging channels to predict the binary responses for both C+ and M+ flares. For works combining both AIA and HMI images, nishizuka2021operational trains a CNN for binary predictions using both C+ and M+ decision rules on full-disk HMI and AIA images, and sun2023tensor proposes a tensor Gaussian process with contraction to combine AIA and HMI images to predict strong flares (M-class) versus weak flares (B-class).
Image Predictors, Continuous Response
Instead of predicting flare classes, another generalization for flare prediction is to directly predict the peak X-ray flux intensity of a flare, now treating the task as a regression problem rather than classification. There is not as much work done along this line as regression is inherently more difficult than classification, but we still provide a brief review here. In muranushi2015ufcorin, the authors use 2D line-of-sight full-disk magnetogram data and then perform a wavelet transform to train a time-series regression model to predict the peak flux intensity of flares. On the other hand, sun2023tensor uses both AIA and HMI imaging data aligned on active regions to perform a tensor regression to predict peak flux intensity. In these works, the models predict a continuous response, but the predictions are then converted into “classification” outputs using flare-class cutoffs in order to use commonly adopted classification metrics such as accuracy and skill scores. However, a work with which they do not bin their continuous predictions is in vandersande2025thes, which investigates key challenges and best practices that should be kept when treating flare prediction as strictly a regression problem.
Vector Predictors, Binary Response
In addition to high-resolution images that pose challenges to model training, many flare-forecasting studies rely on input data vectors that are derived from the summary statistics or data-driven features computed from the images. Space-Weather HMI Active Region Patches (SHARPs)(SHARPs), derived from photospheric magnetic field data taken by the HMI aboard the SDO, provide 25 magnetic and geometric summary parameters. SHARPs are widely used across flare, SEP, and CME forecasting applications. Similar to SHARPs, Space-Weather MDI Active Region Patches (SMARPs)(SMARPs) are derived from maps of the solar surface magnetic field taken by the Michelson Doppler Imager (MDI) aboard the Solar and Heliospheric Observatory (SOHO).
The binary classification of strong flares (such as $\geq$ M-class) against weak flares (such as $\leq C$ or $\leq B$ class) with multi-dimensional time series predictors is the most prevalent in the data-driven solar flare forecasting literature. Previous work on machine learning and deep learning models trained on multidimensional predictors extracted from photospheric magnetic field data (such as SHARPs) shows that carefully engineered vector physical summaries can already achieve competitive performance (bobra2015sola; florios2018forecasting; liu2019predicting; chen2019iden; nishizuka2018deep). Vector predictors are not limited to features derived from images. (verma2026optimal) studies the probability of a strong flare event as an extreme-event prediction using soft X-ray flux time-series vectors.
Vector Predictors, Continuous Response
Flares are classified based on their peak intensity, which is defined to be the peak value of the GOES X-ray flux during the flare; for example, M-class flares have a peak intensity that is at least ${10}^{-5}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$ and less than ${10}^{-4}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$, and X-class flares have a peak intensity that is at least ${10}^{-4}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$ (goes8-15_readme). A typical formulation of the flare forecasting problem is predicting whether an M+ flare will occur in a future 24-hour period (leka2018sola). In this formulation, the outcome is binary. Table 1 provides a justification for an alternative formulation in which the outcome is continuous. The table is based on the SWAN-SF benchmark dataset, which contains various kinds of data, including flare and GOES X-ray flux data, for the period 2010-2018 (angryk2020multivariate). For each date between 1 May 2010 and 29 December 2018, we checked whether an M+ flare occurred and whether the maximum X-ray flux was at or above the M-class threshold of ${10}^{-5}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$. All but four dates either had both an M+ flare and an M+ maximum X-ray flux or had neither. This shows that predicting whether an M+ flare will occur during a 24-hour window is essentially equivalent to predicting whether the maximum X-ray flux during the window will be above the M-class threshold. Predicting the latter event can be done by forecasting the maximum X-ray flux, which is a continuous outcome.
Date Had M+ Flare?
Date Had M+ Maximum X-Ray Flux? Yes No
Yes 434 (17%) 0 (0%)
No 4 (0.15%) 2,146 (83%)
Table 1: The joint distribution of the M+ maximum flux and flare indicators for dates between 1 May 2010 and 29 December 2018, inclusive. The SWAN-SF benchmark dataset described in angryk2020multivariate was used to compute the counts.
Forecasting a continuous outcome has several advantages over forecasting a binary outcome. First, a prediction of a binary outcome is a prediction of whether an event of interest will happen, but not of how severe the event will be if it does happen. A prediction of a continuous outcome is a prediction of both occurrence and severity. Second, the flare class thresholds are arbitrary; there is no substantive difference between the strongest flares in one class and the weakest flares in the next higher class. If ignored, the lack of a clear distinction between adjacent flare classes can make it difficult for a binary forecasting method to perform well, because the method may have trouble correctly predicting flares near a class boundary.
Several works use a continuous outcome instead of a binary outcome. For example, vandersande2025thes models the maximum soft X-ray flux in a 24-hour window using linear regression, random forests, multilayer perceptrons (MLPs), and convolutional neural networks (CNNs). boucheron2015pred employs support vector regression (SVR) trained on magnetic features extracted from MDI images, using a linearly spaced continuous-valued label vector to represent flare size. jiao2020solar develops a mixed LSTM model to forecast the maximum flare intensity, and considers classification models built on top of the regression.
In the sequel, we will describe issues that need to be addressed when using the various kinds of data mentioned previously and suggest ways to address those issues. In Section 2, we discuss response data, covering various data sources, quality, and a systematic comparison. In Section 3, we discuss predictor data, with image-based predictors being discussed in Section 3.1 and vector predictors being discussed in Section 3.2. Finally, in Section 4, we show how predictive performance can be degraded if the issues we raise aren’t addressed.
2 Flare Response: Data Source, Quality, and Processing
2.1 Sources of Data Products for Flare Forecasting
2.1.1 GOES X-ray Flare
[…]
Figure 2: The 1-minute averaged X-Ray flux in log10 scale and the flare events at May 3rd, 2022. The short, vertical lines represent the event time of flares. Not every local maximum in flux will be labeled as one flare event. The flux and the flare events are the Science-Quality data processed by NCEI.
A notable change between the GOES-R (GOES 16-18) and previous GOES data is that the operational (near-real-time) data prior to GOES-16 had scaling factors applied by SWPC to adjust the GOES 8-15 irradiances to match fluxes from GOES-7. Thus, the flare index before GOES-16 was based on the operational irradiances with scaling factors. To get the true irradiances, the scaling factor of 0.7 (for the XRS-B channel) had to be removed (see Section 2.2).
The operational flare data based on the operational 1-minute X-Ray flux is prepared and stored by SWPC. SWPC provides near-real-time and recently observed X-ray flux data, solar and geomagnetic indices and solar event reports created from preliminary reports. Specifically, the SWPC offers Solar Events Lists (SEL) from 1996 to the present, accessible via its public warehouse (see ftp://ftp.swpc.noaa.gov/pub/warehouse/). The SEL provides daily preliminary lists of solar events, such as bright surges on the limb, filament disappearances, optical flares observed in H-alpha, X-ray flares, and others. By selecting the XRA event type, users can access operational X-ray flare event data dating back to 1996 (SEL). We refer to the flare event lists from the SEL as the SWPC-FTP list. The SWPC-FTP list includes flare times, magnitudes, and the assigned NOAA solar active region numbers. Additionally, the warehouse contains the daily SRS (Solar Region Summary) report, which provides detailed information about the solar active regions, including their positions.
Science-Quality XRS data are produced by NOAA’s National Center for Environmental Information (NCEI) and differ from the operational products used at SWPC in that the data have been preprocessed and incorporate retrospective fixes for issues and outages in the operational product. Specifically, NCEI has reprocessed GOES 8-18 operational flare summary data to produce science-quality flare summary data (see Section 2.2). The GOES 1-7 data will be completed in 2026. The Science-Quality flare summary data is based on the 1-minute Science-Quality X-Ray irradiances data. We refer to this flare catalog as the Science-Quality list in this paper. The Science-Quality list does not include the corresponding active region (AR) number for each flare event. However, NCEI provides Science-Quality flare location information for GOES-R (GOES 16-18) data. The flare location on the solar disk is determined based on the measurements from the high flux XRS-B2 quad diode detector and was calibrated and validated with comparisons to flare locations from the Heliophysics Event Registry (goesr_l2_userguide). And this can be used to match to the ARs.
Table 2: Overview of GOES X-ray Flare Sources.
Reference Time Range Satellite Category AR Location
SWPC-FTP 1996/07/31 – present GOES 7–16 Operational Yes No
Sci-Quality 2009/09/01 – 2020/03/04 GOES 8–15 Sci-Quality No No
Sci-Quality 2017/02/09 – present GOES 16–18 Sci-Quality No Yes
SunPy-HEK 1975 – present GOES 1–16 / Yes Yes
2.1.2 AIA Flares and SolarSoft Latest Events
Besides GOES X-Ray flare, AIA flares are solar flares detected and characterized based on extreme ultraviolet (EUV) observations from the Atmospheric Imaging Assembly (AIA lemen2012atmospheric) aboard the Solar Dynamics Observatory (SDO pesnell2012solar). The detection process uses a peak-detection algorithm on binned AIA images, primarily in the 193 Åwavelength, to identify the start, peak, end time, and approximate location of the flare. Basically, the AIA-flares detection algorithm detects flares as intervals of positive derivative and negative derivative around a local maximum (Martens2012).
SolarSoft Latest Events(Freeland2002) is another catalog that regularly detects and collects solar flare events. The catalog is primarily designed to assign spatial locations to GOES flare events by using the best available X-ray/EUV full-disk data. Since 2010, AIA 94åhas been the primary data source, and currently, GOES-R series SUVI 93/94 Åserves as a backup when AIA data is unavailable. It is independent of specific instruments (personal communications with Dr. Samuel Freeland).
2.1.3 HEK Repository
The HEK is a comprehensive metadata system designed to catalog and provide access to solar events, facilitating efficient data retrieval for researchers (hek). HEK serves as an integrated event repository rather than a direct data source. It aggregates flare information from multiple sources. HEK incorporates AIA flare detections from automated feature-finding tools applied to 94 ÅEUV images. It also ingests external sources, including the SWPC operational GOES X-Ray flare list and the SSW catalog.
Additionally, the SunPy package allows users to access solar information through its HEK module. According to the official SunPy document for accessing the GOES flare list, as well as the open source code, users can retrieve the desired SWPC operational flare events from the HEK repository by setting the appropriate event type and observatory.
However, the GOES flare list from the SunPy-HEK module exhibits several discrepancies compared to the flare list available in the SWPC open warehouse, particularly in event counts and Active Region (AR) number information. Between January 1, 1997, and July 21, 2024, 289 flares (1.14%) appear in the SunPy-HEK list but are absent from the SWPC-FTP list, while 817 flares (3.15%) recorded by SWPC are not found in the SunPy-HEK list. Additionally, 9842 flare events have differing AR numbers between the two lists. Figure 3 shows the cumulative flare intensity since January 1, 2010, for both the SunPy-HEK and SWPC-FTP flare lists. A notable disparity appears after 2022, due to 728 flare events recorded by SunPy-HEK that are missing, and more than 3,000 flares are mislabeled with AR number 0 in the SunPy-HEK list. The HEK operation team suggested that the difference in event counts may be due to the double retrieval of the same flare after SWPC revisions (personal communication). Thus, we do not recommend using the SunPy-HEK GOES flare list, especially the GOES-R data, for research purposes.
Figure 3: Cumulative flare intensity since 2010/01/01 for SunPy-HEK and SWPC-FTP flare lists. The slope of the SunPy-HEK without a valid AR number after 2022 is suspiciously steep because more than 3000 flares are mislabeled with AR number 0. The relatively mild difference before 2020 is due to the fact that the different records in the two lists cancel each other out.
2.2 Comparison of Science Quality and Operational Flare Response
NCEI produces and restores science-quality GOES XRS L1b and L2 datasets in netCDF format. The science-quality data have been reprocessed from the start of the mission to the present date and incorporate retrospective fixes for issues and outages in the operational product. The XRS L1b data contains high-cadence soft X-ray irradiance measurements, and the L2 data consists of higher-order products such as irradiance averages, flare event summaries, and flare location products. Science-quality L2 data products are created from science-quality L1b data. Up to October 2025, science-quality flare summary products are available for GOES-R (GOES 16, 17, 18) and GOES 8-15 satellites, dating from 1995/01/03 to the present. 1-minute flux average product, another commonly used L2 product, has been processed to science-quality for GOES 8-18 satellites. This paper focuses mainly on the flare summary product, which contains the most widely used flare response information in flare prediction tasks.
The science-quality data XRS have been reprocessed with many corrections, including smoothness, accuracy, scaling factor consistency, bandpass corrections, and improvement of the data quality flag (goes8-15_readme). The SWPC scaling factor has the most crucial effect on the flare summary product. Science-quality XRS irradiances are provided in physical units of W/m2 (watts per square meter), while the operational data before GOES-16 had SWPC-adjusted scaling factors that adjusted the GOES 8–15 irradiances to match those of GOES 7. Since the flare summary product is based on the L1b data using the flare detection algorithm, the threshold for each flare class increased by 42%, meaning that an individual solar flare before GOES-R needed to be 42% larger than a flare after GOES-R to be given the same level of magnitude. Thus, fewer flares were covered in the operational data set during GOES 8-15. The science-quality data were corrected so that they did not contain the SWPC scaling factors. Figure 4 shows the flare magnitude inconsistency in the operational data. Figure 5 presents the flare number difference related to the SWPC rescaling factors between the operational data and science-quality data.
Figure 4: The flare peak flux ratio of SWPC-FTP list to the NCEI science-quality list through 2020/01/01 - 2024/7/21. Before December 2019, the SWPC applied a rescaling factor such that the ratio is centered at 0.7, while the centered ratio is back to 1 as GOES-16 became the primary operational satellite. The colored rectangle shows the data availability time range for GOES 13-18, which differs from the serving time as the primary satellite providing operational data. GOES-16 became the primary geostationary satellite for its operations in December 2019.
Figure 5: Flare event numbers from NCEI Science-Quality data and SWPC-FTP data. The dashed line represents SWPC-FTP flare records with a nonzero AR number. Prior to December 2019, fewer flares were recorded in the operational data due to the SWPC scaling factor. After the transition to primarily using GOES-16, the number of science-quality flare events is smaller due to the correction.
2.3 Cross-Comparison of Science-Quality, Operational, and SSW Flare Lists
To further look at the differences of these flare lists, we then compared the Science-Quality, SWPC-FTP, and SSW flare lists. The Science-Quality catalog was obtained through the NCEI website. The SWPC-FTP catalog is directly downloaded from the SWPC open warehouse due to the inconsistency between the FTP archive and SunPy-HEK records. The SSW flares were downloaded from their web archive (Freeland2002). Additionally, we downloaded the Science-Quality flare location production from NCEI and augmented the Science-Quality list using the pinpoint location information.
The period of interest covers from 2010-01-01 to 2024-12-31. NCEI provides the processed science-quality data for each instrument separately, and the data availability coverage of one instrument usually does not align with the time range when the instrument is in operational use. For example, GOES-16 reached geostationary orbit on 2015-11-29 and started to provide flare event data since 2017-02-09, while GOES-16 did not become NOAA’s primary source of solar X-ray flux until 2019-12-09. GOES-16 was replaced by GOES-18 as the primary satellite in December 2024.
During this time range, GOES-13, -14, -15, and -16 had become the primary satellites. We compiled the flare summary data from different instruments to get the Science-Quality flare list. Since the flare location product is only available for the GOES-R series and the flare location information is crucial in predictive model training, we include the whole time range of the GOES-16 since 2017-02-19. Before GOES-16, we compiled the science-quality data from different GOES instruments based on primary use status (goes1-15_readme). We show the time range of the compiled science-quality flare data in Table 3.
Table 3: The Science-Quality column indicates the periods used in this study to compile science-quality flare events from various satellites. In general, there have been two operational GOES satellites for each instrument, a primary and a secondary. Primary satellite data should be used over secondary satellite data whenever it is available. There will be gaps in good data for the primary satellites due to eclipses and in-flight calibrations. Satellites are listed in the table only when there are changes. Data courtesy: Kimberly Moreland
Date Science-Quality Primary Secondary
2023-01-04 18
2021-08-24 17
2019-12-09 16 15
2017-12-12 14
2017-08-23 15 13
2017-08-16 14
2017-08-08 13 15
2017-02-07 16
2016-06-09 15 15 13
2016-05-16 15
2016-05-12 14 14 13
2015-06-09 13 13 14
2015-06-09 15 15
2015-05-21 14 14
2015-01-26 13
2012-11-19 15 15 14
2012-10-23 14 14 15
2011-09-01 14
2010-10-28 15 15 None
2010-09-01 15
2009-12-31 14 14 None
During this time range, there are 33215 Science-Quality records, distributed into 139 X-flares and 2497 M-flares. We also downloaded 27586 SWPC-FTP and 33489 SSW flares. We apply minimum data cleaning to these lists, only ensuring the validity of start time, peak time, and flare magnitude features. Table 4 presents the number of flares with different magnitudes in our compiled Science-Quality, SWPC-FTP, and SSW lists.
Table 4: Summary of flare lists from Science-Quality, SWPC-FTP and SSW. The data range is from 2010-01-01 to 2024-12-31.
Source X M C B A Total
Science-Quality 139 2497 19207 11372 0 33,215
SWPC-FTP 126 2193 16049 9120 98 27,586
SSW 124 2194 19791 10023 1357 33,489
For location information, Science-Quality and SSW have pinpoint flare location coordinates, and SWPC-FTP and SSW have NOAA AR numbers information.
We then attempt to match each Science-Quality flare with corresponding events in the SWPC-FTP and SSW catalogs based on temporal characteristics (i.e., flare peak time) and flare intensity. Specifically, we define tolerance thresholds for both peak time and the logarithm (base 10) of the flare intensity. For a given flare, we first filter candidates from the comparison catalog whose absolute difference in peak time falls within the defined time tolerance window. We then further refine the matches by selecting only those flares whose difference in log10 intensity is within the specified threshold. In cases where multiple matches satisfy these criteria, the flare with the closest peak time is selected. Figure 6 shows the fraction of matched flares under various tolerance settings.
Figure 6: Matched flare fraction under different peak-time and intensity tolerances. For each flare, candidate matches are first selected based on peak-time proximity and then filtered by $\log_{10}$-intensity similarity; if multiple candidates remain, the flare with the closest peak time is chosen.
One important consideration is that, before the designation of GOES-16 as the primary satellite, flare intensities reported in the SWPC-FTP catalog were systematically affected by the SWPC scaling factor. This results in reported $\log_{10}$ intensities that are approximately 0.15 lower than their true values. The extent to which the SSW catalog inherits this bias is not explicitly documented. However, our analysis shows that the SSW catalog exhibits a level of similarity to SWPC-FTP that exceeds the similarity observed between any other pair of catalogs. To evaluate this, we first matched flares between the Science-Quality and SSW lists using only peak time with a strict tolerance of zero. For these temporally matched events, we computed the $\log_{10}$ intensity differences across two periods: before 2019-12-07 (prior to GOES-16 availability) and after 2019-12-09 (once GOES-16 data were adopted operationally). Figure 7 shows the resulting histograms. Before GOES-16 became operational, the intensity differences clustered around 0.15, consistent with the expected offset introduced by the SWPC scaling factor. After 2019-12-07, the differences are centered near zero, indicating the removal of this bias. Together with the matching-proportion pattern shown in Figure 6, these results demonstrate that the SSW catalog is not free from the SWPC scaling prior to GOES-16 and, more generally, that SSW aligns more closely with SWPC-FTP than with the Science-Quality catalog. Therefore, the SSW list should be regarded and processed as an operational dataset.
Consequently, for events occurring before 2019-12-09, we adjusted the flare log intensity for SWPC-FTP and SSW by removing the SWPC scaling factor in log scale. In this study, we adopt a peak time tolerance of 15 minutes and a log 10 intensity tolerance of 0.3 as the criteria for identifying corresponding flare events across different catalogs. As shown in Figure 6, when the log intensity tolerance is set over 0.3, the fraction of matched flares remains relatively stable. Figure 8 shows the distribution of Science-Quality flares that fail to match an event in either the operational SWPC-FTP or SSW catalogs. Only 58 flares of class M or higher (2 X-flares) are unmatched, and the overall discrepancy decreases substantially after 2019-12-07, when GOES-R satellites became primarily operational.
Figure 7: Distribution of $\log_{10}$-intensity differences between matched SWPC-FTP and SSW flares. Matching uses a 0-minute peak-time tolerance and a wide log-intensity tolerance, imposing no magnitude constraint. Prior to GOES-16 becoming operational, the intensity differences cluster around 0.15, which is approximately $\log(0.7)$, indicating that the SSW catalog carries the SWPC scaling factor during this period.
Figure 8: Number of solar flare events of each class in years 2010-2024 from science-quality flare list that are not matched to any flare event in the SWPC-FTP or SSW catalogs.
2.4 Science-Quality Flare List Augmentation
We augment the Science-Quality flare list by attaching NOAA Active Region (AR) numbers to individual flares. Location information is essential for flare prediction, as flares arise from localized magnetic activity, and many flare-forecasting studies rely on AR identifiers to link flares to HARP patches and utilize region-specific physical features. However, the NCEI Science-Quality flare summary product does not include NOAA AR numbers. It only provides precise flare locations for the GOES-R series (GOES-16/17/18/19) beginning on 2017-02-09, expressed in four coordinate systems: Stonyhurst/heliographic (longitude, latitude), Carrington coordinates, heliocentric radial coordinates, and helioprojective Cartesian (HPC) $(x,y)$ goes8-15_readme.
For Science-Quality flares that successfully match to a SWPC-FTP or SSW flare, we directly adopt the AR number from the matched operational report. However, approximately 37% of science-quality flares cannot be matched, or are matched to entries associated with an invalid AR number (0). To extend AR coverage for these cases, we use the pinpoint flare locations provided in the GOES-R series together with NOAA Solar Region Summary (SRS) reports.
The SWPC SRS reports, available daily through the public FTP archive, provide detailed information for each numbered active region, including its position in Stonyhurst/heliographic coordinates. For every Science-Quality flare with valid HPC coordinates, we assign an AR number by identifying the nearest active region in the SRS report using the Euclidean distance in HPC coordinates:
$\text{Distance}\;=\;(X_{\text{AR}}-X_{\text{Flare}})^{2}+(Y_{\text{AR}}-Y_{\text{Flare}})^{2}.$
Among the flares occurring during the interval for which precise location information is available, 93% (11,783 events) can be assigned a nearest AR. For those flares that also have a nonzero AR number reported in SWPC-FTP, 70% share the same AR number, and more than 95% differ by $\leq 6$ units. Table LABEL:tab:nearest_ar_distance summarizes the Euclidean distance (in arcseconds) between flare and nearest AR. To ensure reliability, we retain only AR assignments with distance <250<250 arcsec, increasing the number of valid AR labels for approximately 9% of all flares in this period.
For flares occurring before the availability of GOES-R pinpoint locations, AR numbers are assigned only when a valid match exists in either SWPC-FTP or SSW. Eventually, we assign valid AR numbers to 72% of all Science-Quality flares. This augmented science-quality flare list is used in the model-training comparisons presented in Section 4. A complete data-processing pipeline is provided in the Appendix.
Table 5: Summary statistics for Euclidean distance (in arcsec) between Science-Quality flare and the nearest Active Region in the helioprojective Cartesian coordinate. The mean plus one standard deviation is 246.11 arcsec. We set 250 arcsec as the threshold for valid nearest AR choices.
Count mean std. min 50% 75% 90%
6592 151.42 196.13 1.50 102.64 174.01 290
3 Flare Predictor: Data Source, Quality, and Processing
3.1 Imaging Predictors: HMI/SDO, AIA/SDO, and SOHO/MDI
The imaging dataset of interest is the imaging data provided by the Atmospheric Imaging Assembly (AIA) (lemen2012atmospheric) and Helioseismic and Magnetic Imager (HMI) (scherrer2012helioseismic) instruments from the Solar Dynamics Observatory (SDO) since its launch in 2010 (pesnell2012solar). These instruments capture the full-disk images of the Sun in high spatial resolution (4096 $\times 4096$) with high temporal frequency as well. However, the data from HMI and AIA do not align up one-to-one as the AIA images have a larger field of view, resulting in different pixel sizes. There has been much work in processing the SDO instrument data to help facilitate its use in machine learning research due to these kinds of discrepancies. For example, galvez2019machine curates a data set by removing poor quality images, rescaling the images to a lower and more manageable resolution (512 $\times 512$), and applied corrections due to machine degradation and exposure issues.
In particular, the data product that we focus on is tailored specifically for machine-learning based flare prediction research [Jin et al., 2024 in prep], which processes the existing HMI HARP data with the AIA images (in the 8 channels of 94, 131, 171, 193, 211, 304, 35, and 1600 Ångstroms) to make it the same field of view and coordinates. This data set will contain all B, C, M, and X-class flares from 2010-2024, starting 24 hours before the flare peak-time at a temporal cadence of 12 minutes. The flare peak time is described by the Geostationary Operational Environmental Satellite (GOES) which is operated by the United State’s National Oceanic and Atmospheric Administration (NOAA).
The HMI instrument is used to study oscillations and the magnetic field of the photosphere of the Sun’s surface. The image contain 2D photospheric maps of the 3 orthogonal magnetic components. This is in contrast to the MDI instrument from SOHO, which only recorded 2D Line-of-sight imaging data for the magnetic field, which has been popular for previous machine-learning models. In particular, we are interested in the radial field component of the magnetic image $B_{r}$ channel of the HMI images. This orthogonal component of the magnetic field is not only where some of the summary statistics (like SHARPS that we will discuss further in Section 3.2) for the Sun is derived from, but also the magnetic polarity inversion lines (PIL) can be identified from . This is important as the PIL is known to be one of the most important features in triggering flares and other solar weather activity (sun2023tensor).
One issue regarding the solar imaging data from the HMI and AIA that needs to be worried about is with regards to “near-limb” images that are taken. By “near-limb”, we typically mean the images that are taken of active regions that are on the edge, which is a potential data quality issue. As an example, refer to Table 6, which show the mean proportion of missingness for images near the limb and in the Disk center (which we define to have a longitude of above 60 degrees or less than -60 degrees). As we can see, there is a significantly higher proportion of missings when the images are near the limb, with a lot more variability. Thankfully, we have more images on the disk center that are of “higher-quality”,
Region Mean Prop Missing Std. Dev. Total Image Counts
Near Limb 0.114 0.188 424,627
Disk Center 0.00014 0.0035 522,051
Table 6: Mean proportion of missingness for images on the limb versus disk center along with total image counts
For a qualitative perspective of the difference between near limb images and not, Figure 9 is an example of an image of AR11079 of a near limb image. There are clear quality issues with this image, as not only is it blurry and askew, but the specific details and contrast are not very distinct. As such, there is limited information that can possibly be learned by such images.
Figure 9: AR11079 Example Image at 2010-06-13T023100 which is on the limb with a max longitude that is around 80 degrees. It is off-center, blurry, and difficult to make out the specific details.
On the other hand, we have Figure 10 which is of AR11121 when it is near perfectly centered. The quality difference is quite immense, especially as the specific details are quite easy to even visually decipher.
Figure 10: AR11121 Example Image at 2010-11-11 04:36:00 which is a clear, centered image not on the limbs of the instrument.
As such, when dealing with these solar flare images, for quality control is best to only deal with the images that are not on the limb. After such a procedure, we can then curate a list of images for the sake of any modeling or other data tasks.
Another concern with the solar flare imaging data is the sporadic missingness we observe in some videos. For many analyses, we would like a continuous 24-hour sequence of an active region’s evolution prior to flare onset. In practice, such complete coverage is uncommon, because the active region must be identified early enough and remain in view for the entire 24 hours. In addition, rare AIA data gaps can occur which mean that even for regions currently being observed, some time intervals are not recorded, creating “gaps” in otherwise continuous video data.
For a visualization to depict the potential negative effects of this data gap, for the task of X hour again prediction of extreme space weather events like solar flares or SEPs, it may be of interest to use historical imaging data of active regions of 24 hours (or 6 hours or anything of interest) before an eruption as training data for a model. However, if we use a flare list listing all flares from 2010-2024 and their associated active region and look at their available images starting from 1 hour away from peak time to 24 hours away, we can see that there is generally less data available the further away we are trying to forecast. This visualization is in Figure 11.
Figure 11: Temporal coverage of on-disk flare images. (Left) Count of unique flares with available data not on the limbs as a function of nearest hour until flare onset. (Right) Stacked annual counts of unique flares that possess complete (non-limb images) 24-hour data coverage, classified by X-ray intensity (M/X, C, and B classes).
3.2 Vector Predictors: SHARPs
If a machine learning model uses images to predict, then it needs to be trained on a large volume of images in order to predict well. As stated in Section 3.1, HMI images are high-resolution. High-resolution images make for big files and are thus challenging to store in large numbers and to efficiently process in training pipelines. An alternative to using images is to use vectors of summary statistics computed from images; the SHARP dataset is based on this idea. The following description of the dataset draws primarily on (SHARPs).
SHARP stands for Space-weather HMI Active Region Patch. An HMI Active Region Patch, or HARP, is a magnetic structure akin to an active region (HARPs), though a HARP can contain multiple NOAA active regions (SHARPs). A software pipeline automatically detects HARPs in HMI images and tracks them as they traverse the solar disk (HARPs). Every twelve minutes, the pixels within a HARP are used to calculate the vector of sixteen SHARP parameters, each of which describes some aspect of the vector magnetic field of the active region within the HARP. See Table 3 in SHARPs for the definitions and explanations of the SHARPs. For example, the parameter USFLUX measures the total unsigned flux, and the parameter MEANJZD measures the vertical current density. Previous studies have demonstrated that the SHARP parameters have a fair degree of predictive power for forecasting flares (e.g., bobra2015sola; chen2019iden), validating the idea that summary statistics can be an effective alternative to images.
There are two types of HARP products: definitive (DEF-SHARPs) and near-real-time (NRT-SHARPs). See http://jsoc.stanford.edu/HMI/HARPS.html for details. NRT-SHARPs are generated within a few hours of data acquisition and therefore provide more timely access, but their parameter quality is generally lower. Definitive SHARPs become available approximately one week later and incorporate improved calibration and processing. It is important to note that HARP identifiers are not consistent between the definitive and NRT products. The DEF-SHARPs dataset spans an additional two years beyond the availability of NRT-SHARPs (starting on 2010-05-01 versus 2012-09-14), resulting in approximately 2.3 million additional measurement samples that are absent in the NRT set. Even within the overlapping time period, definitive SHARPs include 912 more unique NOAA active regions (ARs) than the NRT product, with only 1043 ARs present in both datasets. Unless otherwise specified, the analyses that follow are based on the definitive SHARPs dataset. We compare these two products in terms of prediction model performance in Section 4.3.
Many other quantities are calculated with the SHARP parameters, such as various longitudes and latitudes. The bits of the QUALITY keyword indicate whether various issues, like eclipses, make the calculated values unreliable; see http://jsoc.stanford.edu/jsocwiki/Lev1qualBits for all the issues that can be flagged. Even if the QUALITY keyword does not indicate any problems, some SHARP parameters can still be missing values; at other times, the entire record, i.e., the SHARP parameters, longitudes and latitudes, QUALITY, etc., may be missing. Keywords that can only be computed when a magnetic field is present can have missing values during “padding intervals”; keywords can also have missing values for HARPs that are faint (HARPs). We do not know if these are the only possible causes of missingness.
We investigated quality issues in the HARP and SHARP data for the period 1 May 2010 to 17 August 2023. The former date is the earliest date for which SHARP data is available (SHARPs). As stated earlier, for any HARP, a new record is created every twelve minutes. The times at which a record can be computed are the start of an hour, 12 minutes after the start, 24 minutes after the start, etc. At each of these times, there may be one or more HARPs in existence, or there may be no HARPs in existence. In the former case, at least one HARP record would be available, and in the latter, no HARP records would be available. For each time in the specified period, we determined whether it lay within the lifetime of some HARP. We found that 526,817 (90%) of the times were within the lifetime of some HARP, and 56,024 (10%) were not within the lifetime of any HARP.
The sequence of times can be partitioned into runs such that within each run, either every time is in some HARP lifetime or no time is in some HARP lifetime. For each run, we computed its start and end times and its duration. These runs are visualized in Figure 12. In the figure, each dot represents one run of times. A run is blue if it consists exclusively of times that lie within the lifetime of some HARP; a run is orange if it only contains times that do not have that property. The first coordinate of a dot is the start time of the corresponding run, while the second coordinate is the run’s duration in days. Orange runs should not be used as there is no HARP data for them and no way to impute it. It seems reasonable to train a model with at least a few months’ worth of data; there are only two blue runs that could be divided into training sets with that much data. Both of those runs correspond to solar maxima.
Figure 12: Each dot in this plot represents one run (see the main text for the definitions of a run and the two types of runs). A dot is colored blue if the corresponding run is of the first type and orange if the run is of the second type. The first coordinate of the dot gives the start time of the run, while the second gives the run length in days. The blue dot in the upper left corner represents the longest run in the data, the run at the top of Table 7. The plot shows that runs starting in the late 2010s tend to be much shorter, with runs of the two types being intermingled. Given the prevalence of runs of the second type in the late 2010s, a reasonable course of action would be to just use the longest run corresponding to the blue dot in the upper left corner to train and evaluate machine learning models.
Table 7 displays the five longest runs that have HARP data at each time. As stated above, only two runs are long enough that they can be carved into multiple reasonably-sized training sets. In the rest of this section, we will only consider times from the longest run, 21:12 on 18 May 2010 to 19:00 on 9 December 2016.
Run Start Time Run End Time Run Length (Days) Run Length (Times)
2010-05-18 21:12:00 2016-12-09 19:00:00 2,396.95 287,635
2021-05-04 01:12:00 2023-08-17 20:00:00 835.78 100,295
2017-06-12 13:00:00 2017-10-08 08:24:00 117.81 14,138
2020-10-07 15:24:00 2021-01-07 19:12:00 92.20 11,065
2016-12-11 07:24:00 2017-03-10 18:00:00 89.44 10,734
Table 7: The time range beginning on 1 May 2010, the first date with SHARP data, can be divided into runs in which either some HARP exists at each time in the run or no HARP exists at any time. A machine learning model cannot be trained or used to generate predictions during a run of the second type. This table shows the start and end times of the five longest runs of the first type. Lengths are given both in days and in times; a count of times is the same as a count of HARP records. The longest run is roughly five-and-a-half years long and contains the maximum of Solar Cycle 24. This run has ample data for training and evaluating machine learning models.
A HARP record can have one or more problems. A record can be missing values for all SHARP parameters; we call such records missing. A record can be missing values for some, but not all SHARP parameters; we call such records incomplete. We flagged records with these problems. It turned out that all SHARP parameters were missing if and only if the entire HARP record was missing. Another problem is that for certain records, some SHARP parameter is infinite. Yet another is that some records are marked as low-quality. We flagged records with these problems. We also flagged records for which part or all of the HARP was more than 68 degrees from the central meridian. Image quality degrades near the limb of the Sun, so the problems for a record could be due to the HARP being far from the central meridian.
As Table 8 shows, most records don’t have any of the problems described above, but a sizable fraction have some problem. Only 0.1% of records have more than one problem. The most important subsets of problematic records consist of records that are only missing (8.1% of records), records that are only low-quality (5.4% of records), and records that are only incomplete (1.3% of records). Records with an infinite value are virtually nonexistent.
Is Record
Missing?
Is Record
Incomplete?
Does Record
Have an $\infty$?
Is Record
Low-Quality?
Number of
Records
Percentage of
Records
No No No No 2,293,123 85%
Yes No No N/A 218,511 8%
No No No Yes 142,805 5%
No Yes No No 33,162 1%
No Yes No Yes 2,683 $\approx$ 0%
No No Yes No 17 $\approx$ 0%
No Yes Yes No 1 $\approx$ 0%
Table 8: A breakdown of the SHARP dataset by data quality problem. A record is missing if the values for all SHARP parameters are missing. A record is incomplete if it has missing values for some SHARP parameters, but not all. A record “has an $\infty$” if the value of some SHARP parameter is $\infty$. A record is low-quality if some bit of the QUALITY keyword is nonzero, which indicates that the record has a quality issue. 85% of records have no problems, and less than 1% have multiple problems. Of those records that do have problems, most are either missing or of low-quality.
4 Comparison of Data Products from a Flare Prediction Perspective
In this section, we compare the SWPC-FTP operational flare list with the NCEI Science-Quality list by evaluating the numerical skill scores of machine learning models trained on definitive or near-real-time SHARP summary vector parameters. Note that we do not exhaust the available machine learning models and do not aim to optimize reported skill scores here, since, as discussed in the Introduction, (i) when properly trained, commonly used machine learning models do not generate significantly different results, and (ii) the skill scores of machine learning models depend heavily on sample selection and sample split schemes. Therefore, we use common practice for the choice of machine learning model and sample split here to demonstrate the differences in the various data products’ effects on the trained model’s performance.
4.1 Data Preparation and Model Setting
To assess how data quality affects flare forecasting performance, we consider two sources of flare labels (the SWPC-FTP list and the Science-Quality list) and two versions of predictor data (HMI near-real-time SHARPs and definitive calibrated SHARPs), see Table 9. We evaluate two representative model families: a deep learning architecture (LSTM) and a conventional statistical model (logistic regression with PCA, i.e., principal component analysis). All models are trained on Solar Cycle 24 (2010-05-01 to 2019-12-31) and tested independently on years 2020–2024. The task is formulated as the prediction of $\geq$ M-class flares in the next 24, 12, 6 hours.
SPWC-FTP Science-Quality
NRT SHARPs O-NRT S-NRT
Definitive SHARPs O-DEF S-DEF
Table 9: Combinations of flare response datasets and SHARP predictor data used in model comparison. For every combination, we train an LSTM-based classifier and a logistic regression model.
We follow the data preparation and sample construction methods described in  (Jiao2020lstm). The input data consists of 24-hour time series extracted from the full time series of SHARP parameters of ARs. To ensure data quality, some sequences are excluded, particularly when the ARs are located near the limb. The criteria for excluding unqualified time sequences are as follows.
  1. 1.
    To minimize projection effects, the longitude of the HARP region must fall within $\pm 70^{\circ}$ of the central meridian.
  2. 2.
    The starting time of the two adjacent time sequences is separated by at least 1 hour.
The response for each input sequence (predictors) is defined as the maximum flare class produced by the corresponding AR within the forecasting window of 24, 12, or 6 hours following the end time of the input sequence. If no flare occurs within this forecasting window, the prediction window is skipped, and the time window is advanced by the stride of 1 hour. Consequently, the target labels can be categorized as ”A,” ”B,” ”C,” ”M,” and ”X”.
In this paper, we focus on a binary classification problem that distinguishes strong(M/X) from Weak (A/B) flares. We exclude the C-class category because its physical significance is ambiguous. C-class flares may be comparable to either small M-class flares or strong B-class flares, making it an unreliable separator boundary and thus vague the training process (chen2019solar). In this study, we adopt two models: an LSTM model with a sigmoid output layer and a logistic regression model. The LSTM-based model is widely used in solar flare forecasting for handling time-series inputs ((chen2019iden), (Liu2019lstm), (Wang2020lstm)) and has been shown to outperform other machine learning models ((Wang2020lstm)). The network consists of a two-layer LSTM with a hidden size of 30 units and dropout of 0.3 between layers, followed by a ReLU activation and a fully connected layer that maps the final (or truncated) hidden representation to a single scalar logit. This logit value is interpreted as the pre-sigmoid score for the probability of a strong flare.
We also include logistic regression due to its simplicity, which in fact results in stability for trained models, and interpretability. Given the $k$-hour forecasting window used to extract SHARP parameters, the input of the LSTM model for each target is a $5k\times 20$ matrix, where each row represents a 20-dimensional vector of SHARP parameters. For the logistic model, we first flatten each input along the time dimension into a $100k\times 1$ vector.
The data are split into training and validation subsets using a 90/10 stratified split to preserve the positive/negative class ratio. To improve generalization and account for sampling variability, we employ a bootstrap ensemble strategy for both LSTM and logistic regression models. Specifically, we draw 30 bootstrap replicates with replacement from the training set. For each replicate, we fit the pipeline that integrates the following steps.
  1. 1.
    Standardization.
  2. 2.
    For the logistic regression model, we then apply principal component analysis (PCA) to reduce dimensionality, retaining 98% variance. For the 24-hr forecasting window, 20 principal components already cover over 98% variance. The projection coefficients for these principal components are derived from the training set.
  3. 3.
    The LSTM model is optimized using the Adam optimizer with a binary cross-entropy loss.
  4. 4.
    The logistic regression model is fitted with $L_{2}$ regularization. The penalty strength is optimized with 5-fold cross-validation and the ROC-AUC scoring objective.
After each bootstrap model is trained, we evaluate probabilistic predictions on the held-out validation set and determine the optimal decision threshold that maximizes the True Skill Statistic (TSS) over a grid of candidate thresholds in $(0,1)$. The optimal threshold and corresponding validation TSS scores are recorded for every bootstrap sample. To properly evaluate the performance of the machine learning algorithms, we split the samples (time sequences of SHARP parameters and their corresponding maximum flare magnitudes) into training and testing sets. Note that splitting by Active Region (AR) or HARP is crucial to prevent information leakage from the training set to the testing set. In this study, all training/testing splits are performed chronologically by HARP, which mimics the operational setting but may not result in the highest skill scores. See wang2020solar for detailed discussions on sample splitting schemes and their impacts on flare prediction results.
It is important to note that varying definitions of positive and negative samples make direct comparisons of results from existing literature challenging, as discussed in chen2024solar. The purpose of this paper is not to propose a universal method for sample construction but rather to compare model performance across different datasets. We also demonstrate numerically that the differences arising from data inconsistency or quality are not affected by the choice of the sample construction method.
4.2 Science-Quality versus Operational “Response”
To quantify the effect of flare-label quality on the prediction models’ skill scores, we compare the LSTM model and the logistic regression model for 6/12/24 forecasting trained using Science-Quality and SWPC-FTP flare list while fixing the SHARPs parameter inputs at definitive SHARPs. Box plots of six commonly adopted skill scores (TSS, HSS, POD, F1, FAR, ACC, see, e.g., chen2019iden for definitions) for the 24-hour forecasting windows are shown in Figure 13 and Figure 14 for the LSTM model and the logistic regression model, respectively. The 6-hour and 12-hour window prediction results exhibit consistent trends as these. The numerical results for all different model setups are given in Tables in Appendix B.
We trained models using data from Solar Cycle 24 (5/1/2010–12/31/2019) and tested them with different time intervals within Solar Cycle 25. The evaluation spans four distinct testing intervals that correspond to different phases of Solar Cycle 25. The 2020–2021 period reflects the early rising phase and is near solar minimum, characterized by extremely low flare occurrence and a highly imbalanced dataset. The 2022 period represents an evolving phase with increasing magnetic complexity and flare frequency. The 2023–2024 interval corresponds to the solar maximum, featuring intense flare activity and a high positive sample ratio (often exceeding 0.8). Finally, the combined 2020–2025 evaluation window averages performance across the full range of solar activity and reflects the overall performance of the model with a balanced test set.
Across the test periods when the solar activity is relatively low, the models trained with SWPC-FTP operational flare list achieve higher predictive skill scores than those trained with Science-Quality labels. This improvement is clearer for logistic models. During the solar minimum (20200101-20211231), the logistic regression model improves from a TSS of 0.15 [0.11,0.18] (Science-Quality) to 0.37 [0.35,0.38] (SWPC-FTP), and the LSTM model improves from 0.42 [0.16,0.59] to 0.57 [0.40,0.69]. Similar margins in performance win appear in other metrics. However, for the LSTM model, the difference is not substantial. However, we do observe more stability with the logistic regression model results (with consistently shorter boxes in Figure 14 as compared to Figure 13) as compared to LSTM results. During the high-activity period (2023–2024) when the positive-sample ratio exceeds 0.85, the Science-Quality-trained models achieve better performance except for the TSS score of the logistic models. Across all the test periods, models trained with SWPC-FTP list get lower FAR scores, demonstrating that models trained with operational flare lists are better at identifying flares without increasing false alarms.
Figure 13: Box plots of six skill scores for different testing year choices (given in the x-axis) in the solar cycle 25 of the M/X versus A/B classification task by the LSTM model trained on definitive SHARPs as inputs with a 24-hour forecasting window. The blue boxes represent the Science-Quality flare list, while the orange boxes represent the SWPC-FP operational flare list. The dashed lines indicate the positive sample proportions in the test set, corresponding to the Y axis on the right-hand-side of each sub-figure. Note that a smaller FAR means better performance. On each box, the black line marks the mean of the 30 independent bootstrap runs, and the triangle marks the median. The lower and upper bounds of the boxes correspond to the first and third quartiles $Q_{1}$ and $Q_{3}$. The upper and lower error bars are at $Q_{3}+1.5(Q_{3}-Q_{1})$ and $Q_{3}-1.5(Q_{3}-Q_{1})$, respectively. The small dots outside of the boxes show the data points outside the error bars (outliers). The mean value and median are calculated including the outliers.
Figure 14: Same as Figure 13 except the model being the logistic regression.
Figure 15 compares the performance of different model configurations during solar minimum and solar maximum, using definitive SHARP inputs. We use the logistic regression model for 6-hour forecasting trained on SWPC-FTP flare lists and definitive SHARPs as the baseline model, and calculate the average performance gain relative to the baseline. The results show that the behavior of the models varies substantially between the two phases of the solar cycle. Here, we focus on the TSS and F1 for comparison.
During solar minimum, LSTM models trained with the Science-Quality flare lists achieve higher TSS and F1 scores than the logistic regression models trained on the same labels. For the SWPC-FTP lists, the LSTM models perform slightly better or about the same as the logistic models. However, during solar maximum, when trained on identical flare lists, the LSTM models perform worse than the logistic models. In addition, the performance differences between LSTM and logistic models are noticeably larger during solar maximum than during solar minimum. This suggests that the more complex flare environment during solar maximum makes model choice more critical and that simpler models as logistic regression can have an advantage.
(a) TSS, Solar Minimum
(b) TSS, Solar Maximum
(c) F1, Solar Minimum
(d) F1, Solar Maximum
Figure 15: Average performance gain relative to the baseline Logreg-O-DEF model for definitive SHARP predictors, evaluated during solar minimum (2020–2022) and solar maximum (2023–2025) testing periods. The baseline model is the logistic regression model for 6-hour forecasting trained on SWPC-FTP flare lists and definitive SHARPs. Each marker represents a model configuration (Logreg-S-DEF, Logreg-O-DEF, LSTM-S-DEF, and LSTM-O-DEF), with color indicating forecasting window length (6 h, 12 h, and 24 h). Marker size is proportional to the standard deviation across 30 bootstrap independent runs. Positive values indicate improvement over the baseline model (black dashed zero line).
4.3 Near-real-time versus Definitive SHARP
We begin by examining the forecasting model performances when using the DEF-SHARPs and NRT-SHARPs datasets. Although the HARP numbers differ between the two products, we find that the conditional flare-response distributions (given the NOAA AR number) are highly similar. For each SHARP observation (every 12 minutes), we match flares from the Science-Quality list using the NOAA active-region number and flare peak time. Observations associated with M- or X-class flares are labeled as positive, and those associated with A- or B-class flares are labeled as negative, consistent with our model formulation. Figure 16 shows the yearly positive rate and positive-sample counts for both DEF-SHARPs and NRT-SHARPs. Despite having fewer valid observations, the NRT-SHARPs product exhibits a positive-rate pattern that closely matches the definitive product across all training years.
Figure 16: Comparison of yearly class imbalance in the SHARPs datasets: near-real-time (NRT-SHARPs, with blue colors) versus definitive (DEF-SHARPs, with orange colors). For each 12-minute SHARPs observation, flares from the Science-Quality list are matched by NOAA active-region number and flare peak time. Observations associated with M- or X-class flares are labeled as positive, while those associated with A- or B-class flares are labeled as negative, consistent with our model definition. Bars show the yearly positive rate (left y-axis), i.e., the proportion of SHARPs matched to M- and X-class flares, for both NRT-SHARPs and DEF-SHARPs. The overlaid line plots (right y-axis) display the annual counts of SHARPs corresponding to positive labels (target = 1) throughout Solar Cycle 24.
Figure 17 presents the average TSS and F1 score gains across eight model configurations. Because the NRT-SHARPs dataset is extremely imbalanced during 2023–2024, we compare performance between solar minimum and the combined evolving + solar maximum periods. During solar minimum, and holding other settings fixed, training on NRT-SHARPs generally improves performance for both logistic and LSTM models. In contrast, during the evolving + solar maximum period, NRT-SHARPs decrease performance, particularly in terms of TSS.
The results also demonstrate that models trained on NRT-SHARPs exhibit greater variability than those trained on DEF-SHARPs, likely due to reduced sample quality and a smaller training sample size. In addition, forecasting performance across different forecasting-window lengths is more consistent when models are trained on DEF-SHARPs.
(a) TSS, Solar Minimum
(b) TSS, Evolving + Solar Maximum
(c) F1, Solar Minimum
(d) F1, Evolving + Solar Maximum
Figure 17: Average performance gain relative to the baseline Logreg-O-NRT model for Near-Real-Time SHARP predictors, evaluated during solar minimum (2020–2022) and Evolving + solar maximum (2022–2025) testing periods. The baseline model is the logistic regression model for 6-hour forecasting trained on SWPC-FTP flare lists and NRT-SHARPs. Each marker represents a model configuration, with color indicating forecasting window length (6 h, 12 h, and 24 h). Marker size is proportional to the standard deviation across 30 bootstrap independent runs. Positive values indicate improvement over the baseline model (black dashed zero line).
5 Conclusion
In this paper, we systematically review and sort the most commonly used data products in solar flare forecasting with machine-learning methods. For the flare-response labels, we cross-evaluate the Science-Quality list from NCEI, the operational SWPC-FTP list, and the SSW catalog. We also construct an augmented Science-Quality flare list with assigned NOAA active-region numbers and develop a reproducible processing pipeline for generating this enhanced dataset. For predictor data, we describe the imaging products from AIA/SDO and HMI/SDO and analyze the quality issues that arise near the solar limb. We further examine completeness, missingness, and other quality concerns in the HARP and SHARP parameter datasets.
In addition, we compare the predictive performance of LSTM-based models and logistic-regression baselines across multiple combinations of flare lists (Science-Quality and SWPC-FTP) and SHARP data (DEF-SHARPs and NRT-SHARPs). Our results highlight several important conclusions:
  1. 1.
    The Science-Quality flare list does not yield notable improvements in forecasting skill, particularly during solar minimum. This may be explained by information loss: only about 72% of Science-Quality flares can be assigned a valid active region number and thus incorporated into model training.
  2. 2.
    NRT-SHARPs introduce substantially greater variability in model skill scores, likely resulting from reduced data quality and a smaller effective training sample.
  3. 3.
    The impact of data products is strongly solar-cycle dependent. Holding other factors fixed, the operational SWPC-FTP labels improve model performance during solar minimum but degrade performance during solar maximum; NRT-SHARPs exhibit a similar pattern.
  4. 4.
    Although the more sophisticated LSTM models do not consistently outperform logistic regression, they produce more stable results across forecasting windows and data product combinations, indicating robustness to label and predictor variations.
By clarifying the strengths and limitations of each product and quantifying their effects on predictive performance, we hope to help researchers in selecting appropriate datasets and to improve the comparability and interpretability of data-driven flare-forecasting studies.
Acknowledgments
The authors acknowledge NASA for funding the CLEAR Space Weather Center of Excellence (grant number 80NSSC24K0062).
Appendix A Quiet Period Analysis of Active Regions
In this section, we are trying to clearly identify if there are any distributional differences when an active region has an ongoing flare, versus when it is during a “quiet period”. As an initial approach, we are defining the ”quiet period” of an active region to be any period of time that is 24 hours removed from any flare activity (i.e. more than 24 hours before the flare start time and after the flare ends). We are measuring these distributional differences by looking at the marginal distribution of all the SHARPs depending on if it’s a quiet period or a flaring period (which is defined to be any time that is not a quiet period).
After running through all the flares from 2010 to 2024 and categorizing the periods of times of all active regions that existed during this interval as quiet or flaring, we analyzed the marginal distriubtion of the SHARPs during this time. These are visualized here in Figure 18
Figure 18: Marginal Distribution of SHARPs over different periods aggregated across Active Regions.
The quiet (blue) line depicts the SHARP parameters during the quiet period of an AR. The green line depicts the SHARP parameters of active regions that are available in the combined SHARPs data but do not have any flares associated with them in the GOES flare list. For the active regions present in the GOES flare list, we also visualized the distribution of the SHARPs during B-flares to investigate if B-flares could serve as a “‘surrogate” for quiet time. The red curve depicts the exact time interval of a B-flare as described by the GOES flare list, and the purple line expands the interval by 24 hours before and after the start and end times, respectively.
This plot shows that the distribution of a few SHARPs differs depending on whether it is during a quiet period or when there is activity. In particular, the distribution of MEANGAM, MEANSHR, and SHGRT45 suggests that there are indeed differences in the behavior of active regions depending on what time period they are in. To better generalize the behavior of active regions, it may be beneficial to incorporate quiet time data of active regions as well.
Additionally, Figure 19 depicts the class of all flares in the GOES flare list starting from 1996, which encapsulates two solar cycles. During the peak of the solar cycle, we can see that a percentage of B-flares is missing. This indicates that the flare list is not necessarily capturing all B-flares, as they could be obscured by the stronger flares that occur more frequently during the solar cycle. As such, the quiet time imaging data during these solar cycle peaks can also serve as a “representation” of the B-flares during this time.
Figure 19: Marginal Distribution of SHARPs over different periods aggregated across Active Regions.
Additionally, the two dashed lines represent the potential shift that is necessary for the GOES flare labels. As discussed in Section 2, there was an erroneous scaling factor being applied to flares measured prior to the GOES-16 satellite. If we adjust for this scaling, all the flares from before 2019 will be shifted up the length from the “black” line” to the red line. In this scenario, we will have systematically missing low-intensity B-flares in our data, so having the quiet data during this time period can help mitigate any potential negative effects of this.
Appendix B Comparison of Numerical Results
Table 10: Summary of LSTM model performance for 24-hr forecasting using definitive SHARP parameters with Science-Quality flares (S-DEF) and SWPC-FTP flares (O-DEF) across three independent testing periods. Each cell reports the mean and range (minimum–maximum) of bootstrap ensemble results. The rows correspond to varying levels of solar activity: “Min” (2020–2021), “Evolving” (2022), “Max” (2023–2024), and “All” (2020-2024)
Period Model TSS HSS POD F1 FAR ACC
Min LSTM-S-DEF 0.42 [0.16, 0.59] 0.09 [0.03, 0.13] 0.75 [0.45, 0.91] 0.15 [0.10, 0.19] 0.32 [0.28, 0.39] 0.68 [0.62, 0.72]
LSTM-O-DEF 0.57 [0.40, 0.69] 0.19 [0.14, 0.24] 0.82 [0.64, 0.96] 0.26 [0.21, 0.30] 0.25 [0.21, 0.30] 0.75 [0.71, 0.79]
Evolving LSTM-S-DEF 0.59 [0.50, 0.68] 0.60 [0.51, 0.69] 0.71 [0.62, 0.78] 0.73 [0.66, 0.79] 0.12 [0.07, 0.20] 0.82 [0.78, 0.87]
LSTM-O-DEF 0.58 [0.53, 0.66] 0.61 [0.56, 0.69] 0.64 [0.58, 0.72] 0.73 [0.70, 0.79] 0.06 [0.03, 0.10] 0.82 [0.80, 0.86]
Max LSTM-S-DEF 0.44 [0.39, 0.50] 0.09 [0.07, 0.11] 0.46 [0.40, 0.52] 0.63 [0.57, 0.69] 0.01 [0.00, 0.04] 0.49 [0.43, 0.55]
LSTM-O-DEF 0.38 [0.31, 0.44] 0.05 [0.04, 0.07] 0.38 [0.31, 0.44] 0.55 [0.48, 0.61] 0.00 [0.00, 0.00] 0.41 [0.34, 0.47]
All LSTM-S-DEF 0.57 [0.49, 0.62] 0.57 [0.50, 0.63] 0.63 [0.56, 0.71] 0.74 [0.69, 0.79] 0.06 [0.04, 0.10] 0.79 [0.76, 0.82]
LSTM-O-DEF 0.51 [0.46, 0.58] 0.52 [0.46, 0.58] 0.55 [0.49, 0.61] 0.69 [0.65, 0.74] 0.03 [0.01, 0.05] 0.76 [0.73, 0.79]
Table 11: Summary of LSTM model performance for 12-hr forecasting using definitive SHARP parameters with Science-Quality flares (S-DEF) and SWPC-FTP flares (O-DEF) across three independent testing periods. Each cell reports the mean and range (minimum–maximum) of bootstrap ensemble results. The rows correspond to varying levels of solar activity: “Min” (2020–2021), “Evolving” (2022), “Max” (2023–2024), and “All” (2020-2024)
Period Model TSS HSS POD F1 FAR ACC
2020–2022 LSTM-S-DEF 0.32 [0.07, 0.45] 0.05 [0.01, 0.07] 0.65 [0.38, 0.82] 0.10 [0.06, 0.12] 0.33 [0.30, 0.37] 0.67 [0.63, 0.70]
LSTM-O-DEF 0.50 [0.34, 0.64] 0.12 [0.08, 0.18] 0.77 [0.61, 0.92] 0.18 [0.15, 0.23] 0.27 [0.23, 0.31] 0.73 [0.70, 0.78]
2022–2023 LSTM-S-DEF 0.59 [0.53, 0.70] 0.58 [0.52, 0.67] 0.72 [0.63, 0.83] 0.70 [0.66, 0.77] 0.13 [0.07, 0.17] 0.83 [0.80, 0.87]
LSTM-O-DEF 0.59 [0.52, 0.63] 0.63 [0.54, 0.68] 0.65 [0.61, 0.71] 0.74 [0.68, 0.77] 0.06 [0.03, 0.11] 0.84 [0.80, 0.87]
2023–2025 LSTM-S-DEF 0.44 [0.38, 0.50] 0.10 [0.08, 0.12] 0.45 [0.39, 0.52] 0.62 [0.56, 0.68] 0.01 [0.00, 0.02] 0.48 [0.43, 0.55]
LSTM-O-DEF 0.37 [0.30, 0.45] 0.06 [0.04, 0.08] 0.37 [0.31, 0.45] 0.54 [0.47, 0.62] 0.00 [0.00, 0.02] 0.41 [0.34, 0.48]
2020–2025 LSTM-S-DEF 0.58 [0.50, 0.66] 0.60 [0.52, 0.67] 0.65 [0.56, 0.73] 0.75 [0.68, 0.80] 0.06 [0.04, 0.08] 0.81 [0.77, 0.84]
LSTM-O-DEF 0.53 [0.47, 0.58] 0.55 [0.49, 0.60] 0.57 [0.50, 0.63] 0.70 [0.65, 0.74] 0.04 [0.02, 0.07] 0.79 [0.76, 0.81]
Table 12: Summary of LSTM model performance for 6-hr forecasting using definitive SHARP parameters with Science-Quality flares (S-DEF) and SWPC-FTP flares (O-DEF) across three independent testing periods. Each cell reports the mean and range (minimum–maximum) of bootstrap ensemble results. The rows correspond to varying levels of solar activity: “Min” (2020–2021), “Evolving” (2022), “Max” (2023–2024), and “All” (2020-2024)
Period Model TSS HSS POD F1 FAR ACC
2020–2022 LSTM-S-DEF 0.33 [0.03, 0.48] 0.04 [0.00, 0.06] 0.66 [0.37, 0.85] 0.08 [0.05, 0.10] 0.33 [0.27, 0.37] 0.67 [0.63, 0.72]
LSTM-O-DEF 0.49 [0.30, 0.60] 0.09 [0.05, 0.13] 0.77 [0.60, 0.93] 0.14 [0.11, 0.17] 0.29 [0.24, 0.34] 0.71 [0.66, 0.77]
2022–2023 LSTM-S-DEF 0.57 [0.52, 0.63] 0.55 [0.49, 0.61] 0.73 [0.65, 0.79] 0.67 [0.63, 0.72] 0.15 [0.10, 0.19] 0.82 [0.79, 0.85]
LSTM-O-DEF 0.60 [0.52, 0.66] 0.62 [0.55, 0.69] 0.69 [0.60, 0.79] 0.72 [0.67, 0.77] 0.09 [0.05, 0.16] 0.84 [0.82, 0.88]
2023–2025 LSTM-S-DEF 0.43 [0.34, 0.47] 0.11 [0.08, 0.13] 0.43 [0.35, 0.48] 0.61 [0.52, 0.65] 0.01 [0.00, 0.02] 0.48 [0.40, 0.52]
LSTM-O-DEF 0.36 [0.25, 0.48] 0.07 [0.04, 0.10] 0.36 [0.26, 0.48] 0.53 [0.41, 0.65] 0.00 [0.00, 0.04] 0.40 [0.31, 0.51]
2020–2025 LSTM-S-DEF 0.58 [0.48, 0.65] 0.60 [0.51, 0.67] 0.66 [0.56, 0.72] 0.74 [0.67, 0.79] 0.08 [0.05, 0.11] 0.82 [0.78, 0.84]
LSTM-O-DEF 0.53 [0.49, 0.64] 0.56 [0.51, 0.65] 0.60 [0.52, 0.74] 0.71 [0.66, 0.79] 0.06 [0.03, 0.10] 0.80 [0.78, 0.84]
Table 13: Summary of Logistic model performance for 24-hr forecasting using definitive SHARP parameters with Science-Quality flares (S-DEF) and SWPC-FTP flares (O-DEF) across three independent testing periods. Each cell reports the mean and range (minimum–maximum) of bootstrap ensemble results. The rows correspond to varying levels of solar activity: “Min” (2020–2021), “Evolving” (2022), “Max” (2023–2024), and “All” (2020-2024)
Period Model TSS HSS POD F1 FAR ACC
Min Logistic-S-DEF 0.15 [0.11, 0.18] 0.07 [0.06, 0.09] 0.27 [0.23, 0.30] 0.13 [0.11, 0.14] 0.12 [0.11, 0.13] 0.86 [0.84, 0.87]
Logistic-O-DEF 0.37 [0.35, 0.38] 0.24 [0.23, 0.25] 0.46 [0.43, 0.48] 0.29 [0.29, 0.30] 0.09 [0.09, 0.10] 0.88 [0.88, 0.89]
Evolving Logistic-S-DEF 0.62 [0.61, 0.64] 0.65 [0.64, 0.67] 0.69 [0.69, 0.71] 0.75 [0.75, 0.77] 0.07 [0.07, 0.07] 0.85 [0.85, 0.86]
Logistic-O-DEF 0.68 [0.67, 0.69] 0.71 [0.70, 0.71] 0.73 [0.72, 0.74] 0.81 [0.80, 0.81] 0.05 [0.05, 0.06] 0.87 [0.86, 0.87]
Max Logistic-S-DEF 0.69 [0.67, 0.70] 0.27 [0.26, 0.28] 0.79 [0.78, 0.80] 0.88 [0.87, 0.89] 0.10 [0.09, 0.13] 0.80 [0.79, 0.81]
Logistic-O-DEF 0.74 [0.72, 0.76] 0.21 [0.21, 0.22] 0.76 [0.76, 0.77] 0.87 [0.86, 0.87] 0.02 [0.01, 0.05] 0.77 [0.77, 0.78]
All Logistic-S-DEF 0.66 [0.65, 0.67] 0.66 [0.65, 0.67] 0.76 [0.75, 0.77] 0.81 [0.81, 0.82] 0.10 [0.10, 0.11] 0.83 [0.83, 0.84]
Logistic-O-DEF 0.67 [0.66, 0.67] 0.67 [0.66, 0.67] 0.75 [0.74, 0.75] 0.82 [0.81, 0.82] 0.08 [0.07, 0.08] 0.83 [0.83, 0.84]
Table 14: Summary of Logistic model performance for 12-hr forecasting using definitive SHARP parameters with Science-Quality flares (S-DEF) and SWPC-FTP flares (O-DEF) across three independent testing periods. Each cell reports the mean and range (minimum–maximum) of bootstrap ensemble results. The rows correspond to varying levels of solar activity: “Min” (2020–2021), “Evolving” (2022), “Max” (2023–2024), and “All” (2020-2024)
Period Model TSS HSS POD F1 FAR ACC
Min Logistic-S-DEF 0.12 [0.05, 0.17] 0.04 [0.02, 0.06] 0.25 [0.18, 0.30] 0.09 [0.06, 0.11] 0.13 [0.12, 0.13] 0.86 [0.85, 0.86]
Logistic-O-DEF 0.28 [0.22, 0.38] 0.14 [0.11, 0.19] 0.39 [0.32, 0.49] 0.19 [0.17, 0.24] 0.11 [0.10, 0.11] 0.87 [0.87, 0.88]
Evolving Logistic-S-DEF 0.58 [0.56, 0.61] 0.60 [0.58, 0.64] 0.65 [0.63, 0.68] 0.71 [0.69, 0.73] 0.08 [0.07, 0.09] 0.85 [0.84, 0.86]
Logistic-O-DEF 0.64 [0.62, 0.66] 0.68 [0.66, 0.70] 0.69 [0.66, 0.71] 0.77 [0.76, 0.79] 0.04 [0.04, 0.05] 0.87 [0.86, 0.87]
Max Logistic-S-DEF 0.66 [0.65, 0.67] 0.28 [0.28, 0.29] 0.78 [0.78, 0.79] 0.87 [0.87, 0.88] 0.12 [0.11, 0.13] 0.79 [0.78, 0.80]
Logistic-O-DEF 0.71 [0.67, 0.73] 0.23 [0.22, 0.24] 0.76 [0.75, 0.77] 0.86 [0.86, 0.87] 0.05 [0.03, 0.08] 0.77 [0.76, 0.78]
All Logistic-S-DEF 0.64 [0.63, 0.65] 0.64 [0.64, 0.65] 0.75 [0.74, 0.76] 0.79 [0.79, 0.80] 0.11 [0.10, 0.12] 0.83 [0.82, 0.83]
Logistic-O-DEF 0.65 [0.64, 0.65] 0.66 [0.65, 0.66] 0.73 [0.72, 0.74] 0.80 [0.79, 0.80] 0.09 [0.08, 0.09] 0.83 [0.83, 0.84]
Table 15: Summary of Logistic model performance for 6-hr forecasting using definitive SHARP parameters with Science-Quality flares (S-DEF) and SWPC-FTP flares (O-DEF) across three independent testing periods. Each cell reports the mean and range (minimum–maximum) of bootstrap ensemble results. The rows correspond to varying levels of solar activity: “Min” (2020–2021), “Evolving” (2022), “Max” (2023–2024), and “All” (2020-2024)
Period Model TSS HSS POD F1 FAR ACC
Min Logistic-S-DEF 0.02 [-0.10, 0.20] 0.01 [-0.03, 0.06] 0.16 [0.04, 0.33] 0.05 [0.01, 0.09] 0.14 [0.13, 0.15] 0.84 [0.83, 0.86]
Logistic-O-DEF 0.25 [0.13, 0.27] 0.10 [0.06, 0.11] 0.37 [0.24, 0.38] 0.14 [0.10, 0.16] 0.12 [0.10, 0.13] 0.87 [0.85, 0.88]
Evolving Logistic-S-DEF 0.52 [0.48, 0.57] 0.56 [0.53, 0.61] 0.59 [0.54, 0.62] 0.66 [0.63, 0.70] 0.07 [0.06, 0.08] 0.84 [0.84, 0.86]
Logistic-O-DEF 0.55 [0.52, 0.59] 0.60 [0.57, 0.64] 0.59 [0.55, 0.64] 0.70 [0.67, 0.73] 0.04 [0.03, 0.05] 0.85 [0.84, 0.86]
Max Logistic-S-DEF 0.67 [0.63, 0.71] 0.33 [0.31, 0.34] 0.79 [0.78, 0.80] 0.88 [0.87, 0.88] 0.12 [0.07, 0.15] 0.80 [0.79, 0.80]
Logistic-O-DEF 0.68 [0.67, 0.72] 0.24 [0.23, 0.26] 0.74 [0.72, 0.76] 0.85 [0.84, 0.86] 0.06 [0.03, 0.09] 0.75 [0.73, 0.77]
All Logistic-S-DEF 0.63 [0.62, 0.64] 0.64 [0.63, 0.65] 0.75 [0.73, 0.76] 0.78 [0.77, 0.78] 0.12 [0.11, 0.13] 0.83 [0.82, 0.83]
Logistic-O-DEF 0.61 [0.59, 0.63] 0.63 [0.61, 0.64] 0.70 [0.68, 0.72] 0.76 [0.75, 0.77] 0.09 [0.08, 0.10] 0.82 [0.82, 0.83]
Appendix C Data Processing Pipeline
Figure 20: A diagram illustrating the data processing pipeline.

### v2

@@
-Data Quality Issues in Flare Prediction using Machine Learning Models
+Defects and Inconsistencies in Solar Flare Data Sources: Implications for Machine Learning Forecasting
 Abstract
-Machine learning models for forecasting solar flares have been trained and tested using a variety of data sources, such as Space Weather Prediction Center (SWPC) operational and science-quality data.
-Typically, data from these sources is minimally processed before being used to train and validate a forecasting model.
-However, predictive performance can be impaired if defects in and inconsistencies between these data sources are ignored.
-For a number of commonly used data sources, together with softwares that query and then output processed data, we identify their respective defects and inconsistencies, quantify their extent, and show how they can affect the predictions produced by data-driven machine learning forecasting models.
+Machine learning models for forecasting solar flares have been trained and evaluated using a variety of data sources, including Space Weather Prediction Center (SWPC) operational and science-quality data.
+Typically, data from these sources are minimally processed before being used to train and validate a forecasting model.
+However, predictive performance can be affected if defects and inconsistencies between these data sources are ignored.
+For a number of commonly used data sources, together with software that queries and outputs processed data, we identify their respective defects and inconsistencies, quantify their extent, and show how they can affect the predictions produced by data-driven machine-learning forecasting models.
 We also outline procedures for fixing these issues or at least mitigating their impacts.
-Finally, based on our thorough comparisons of the impacts of data sources on the trained forecasting model in terms of predictive skill scores, we offer recommendations for the use of different data products in operational forecasting.
+Finally, based on thorough comparisons of the effects of data sources on the trained forecasting model’s predictive skill scores, we offer recommendations for the use of different data products in operational forecasting.
 Keywords: solar flares, data quality, machine learning, operational forecasting.
 1 Introduction
-In recent years, machine learning methods have been applied with great success to a variety of space weather forecasting problems, such as solar eruption forecasting (Leka:2018b; leka2019comparison; Camporeale:2019a; chen2019iden; kasapis2022interpretable; whitman2023review; chen2025decent), geomagnetic index prediction (iong2022new; iong2024sparse; chen2025geodgp), and the prediction of terrestrial impacts of space weather events (wang2023forecast; sun2025matrix).
-These methods have also been used for solar flare forecasting, for which they have improved the state of the art (jiao2020solar; sun2021improved; sun2022predicting; sun2023tensor; chen2024solar).
-A machine learning method predicts the value of a response variable from the values of one or more predictor variables.
-In solar flare forecasting in particular, the response is typically the flare class or maximum soft X-ray flux during a flare; the predictors may include remove sensing images such as those from the Solar Dynamics Observatory (SDO) (Boerner:2012a; lemen2012atmospheric; pesnell2012; woods2012; sun2023tensor) or extracted features like the Space eather HMI Active Region Patch (SHARP) parameters (SHARPs).
-The method is trained to generate predictions via minimization of a loss function that measures the degree to which predicted outcomes deviate from observed outcomes; during training, the mapping from predictors to predictions is manipulated (e.g., by gradient descent) to decrease the loss.
-Common choices of loss function in the flare forecasting literature include the cross-entropy loss (for binary classification of flares as strong or weak, e.g., chen2019iden), sum of squares loss (for prediction of the logarithm of the peak flux, e.g., jiao2020solar), and the negative log-likelihood (in either classification or regression tasks with statistical models, e.g., sun2023tensor).
-State-of-the-art binary classification methods for forecasting flares include discriminant analysis (leka2018nwra), logistic regression, support vector machines (bobra2015sola), random forests, gradient boosting machines, long-short term memory (LSTM) neural networks (liu2019predicting), and other more advanced network structures.
-The literature shows that the performance of machine learning methods is more constrained by the train/validation/test design (random split versus chronological split, for example) and sample size (collecting non-overlapping flare events versus rolling window definition of maximum flaring events) than by the particular machine learning method adopted, due to solar flare data sparsity; see Wang2020lstm; sun2022predicting; chen2024solar for more discussion.
+Machine learning methods have been applied with great success to a variety of space weather forecasting problems, such as solar eruption forecasting (e.g.
+Leka and Barnes, 2018b; Leka et al., 2019; Chen et al., 2019a; Kasapis et al., 2022; Whitman et al., 2023; Chen et al., 2025a), geomagnetic index prediction (e.g.
+Iong et al., 2022; Hu et al., 2023; Nair et al., 2023; Iong et al., 2024; Chen et al., 2025b), and the prediction of terrestrial impacts of space weather events (e.g.
+Wang et al., 2019; Telloni et al., 2023; Wang et al., 2023; Elfiky et al., 2025; Sun et al., 2025); see Camporeale et al.
+(2018) and Camporeale (2019) for reviews of methods and applications.
+In particular, machine learning methods have been used for solar flare forecasting, showing superior performance as compared with state-of-the-art physics-based models (e.g.
+Florios et al., 2018; Liu et al., 2019b; Sun et al., 2021; Georgoulis et al., 2021; Nishizuka et al., 2021; Sun et al., 2022; Zheng et al., 2023; Pandey et al., 2023).
+A machine learning forecasting method typically predicts the value of a response variable from one or more predictor variables.
+In solar flare forecasting, in particular, the response is typically the flare class (Nishizuka et al., 2018) or maximum soft X-ray flux during a flare (Yi et al., 2020); the predictors may include remote sensing images such as those from the Solar Dynamics Observatory (SDO) (Boerner et al., 2012; Lemen et al., 2012; Pesnell et al., 2012b; Woods et al., 2012; Sun et al., 2023) or extracted features like the Space eather HMI Active Region Patch (SHARP) parameters (Bobra et al., 2014).
+The method is trained to generate predictions by minimizing a loss function that quantifies the extent to which predicted outcomes deviate from observed outcomes.
+During training, the mapping from predictors to predictions is optimized (e.g., via gradient-based methods) to minimize the loss.
+Common choices of loss function in the flare forecasting literature include the cross-entropy loss (for binary classification of flares as strong or weak, e.g., Chen et al.
+(2019a)), sum of squares loss (for prediction of the logarithm of the peak flux, e.g., Jiao et al.
+(2020b)), and the negative log-likelihood (in either classification or regression tasks with statistical models, e.g., Sun et al.
+(2023)).
+State-of-the-art binary classification methods for forecasting flares include discriminant analysis (Leka et al., 2018), logistic regression (Korsós et al., 2021), support vector machines (Bobra and Couvidat, 2015), random forests (Liu et al., 2017), gradient-boosting machines (Cinto et al., 2020), long short-term memory (LSTM) neural networks (Liu et al., 2019b), and other more advanced network architectures (e.g.
+Zheng et al., 2019; Sun et al., 2022; Pandey et al., 2024).
+The literature shows that the performance of machine learning methods is more constrained by the train, validation, and test design (random split versus chronological split, for example) and sample size (collecting non-overlapping flare events versus rolling window definition of maximum flaring events) than by the particular machine learning method adopted, due to solar flare data sparsity, especially strong M/X class flares; see Wang et al.
+(2020b); Sun et al.
+(2022); Chen et al.
+(2024) for more discussions.
 Therefore, in this paper, instead of exhaustively evaluating all available machine learning methods, we select representative ones (to be detailed later) and focus on using them to calibrate the impact of the data sources (from different databases) and quality (scientific or operational) we use for training the machine learning model.
@@
 The images and plots display data from around the time of an X8.7-class flare that occurred on 14 May 2024 in NOAA active region 13664 (HARP 11149).
-Different pairs of predictor and outcome types have been used in solar flare forecasting literature.
-In Figure 1, we provide an overview of how predictions are produced by a machine learning method from different kinds of predictor data for both classification and regression in solar flare forecasting.
-Predictors are commonly computed from either images of active regions or summary statistics on those regions.
-The upper left part of the schematic depicts predictors being computed from Helioseismic and Magnetic Imager (HMI) and (Atmospheric Imaging Assembly) AIA images, while the lower left part depicts predictors being computed from SHARP parameters, which are summary statistics calculated from HMI images.
-If the method performs classification, then the outcome is the binary variable of whether a flare will occur; if the method performs regression, then the outcome is a number, like the future peak soft X-ray flux.
+1.1 Predictor-Outcome Combinations in Solar Flare Forecasting
+Different pairs of predictor and outcome types have been used in the solar flare forecasting literature.
+In Figure 1, we provide an overview of how predictions are produced by a machine learning method from different types of predictor data for both classification and regression in solar flare forecasting.
+Predictors are commonly computed from either the images of active regions or summary statistics on those regions.
+The upper left part of the schematic depicts predictors being computed from Helioseismic and Magnetic Imager (HMI) and (Atmospheric Imaging Assembly) AIA images, while the lower left part depicts predictors being computed from SHARP parameters, which are summary statistics calculated from HMI measurements.
+If the method performs classification, the outcome is a binary variable indicating whether a flare will occur; if the method performs regression, the outcome is a numeric variable, such as the future peak soft X-ray flux.
 The upper and lower right panels show comparisons of actual outcomes to mock predictions for classification and regression, respectively.
-The images and plots display data from around the time of an X8.7-class flare that occurred on 14 May 2024 in NOAA active region 13664 (HARP 11149).
+The images and plots present data from around the time of an X8.7-class flare on 14 May 2024 in NOAA active region 13664 (HARP 11149).
 Here we briefly discuss some relevant works corresponding to the parts illustrated in Figure 1.
 Detailed numerical results will be presented in Sections 2 and 4.
-Image Predictors, Binary Response
+Here, we briefly review the different predictor-outcome combinations in solar flare forecasting literature with machine learning methods.
+1.1.1 Image Predictors and Binary Responses
 Image predictors typically consist of multi-channel AIA and HMI images of the Sun.
-Full-disk images of the Sun at their native resolution of $4096\times 4096$ for each channel are usually too computationally and memory-intensive to be used as inputs to machine learning models, even when performing batch learning.
+Full-disk images of the Sun at their native resolution of $4096\times 4096$ for each channel are usually too computationally and memory-intensive to serve as inputs to machine learning models, even when using batch learning.
 However, resizing these images through down-sampling or other interpolation methods can degrade spatial perception and erase meaningful, fine-grained details that are useful for flare prediction.
-On the other hand, multi-channel images focused on active regions [Jin et al., 2024, in prep] must not only deal with potential concurrent temporal effects (for active regions whose timelines overlap) but also with varying image resolutions, as the sizes and shapes of active regions differ.
-To address this, works such as Pandey2024 perform a rectangular crop that maximizes unsigned flux (a SHARP parameter).
-However, other work, such as sun2023tensor, uses a similar but different approach by performing a crop centered on the polarity inversion line (PIL).
-PILs are typically the main source of flare activity, as shown in many works e.g.
-Zirin:1993; Green:2018; sun2021improved, so such crops can not only keep predictive power high but also allow for succinct comparisons of features (such as topological and spatial features) the model detects with the derived PIL for potential interpretability results.
-Using these multi-channel images of active regions is quite common for flare prediction.
-In the most basic case, we can develop and train machine learning models to predict whether or not a flare will be “strong” within some time horizon.
-In this case, the response is binary, with “strong” typically meaning C-class or higher (C+) or M-class or higher (M+).
-While this kind of forecasting may seem simpler than the kinds described below, it can still be difficult.
-For example, if only M+ flares are considered strong, forecasting can be difficult because of class imbalance and because of the lack of a clear distinction between flares just below and just above the M-class boundary.
-Instead of just predicting whether a flare will be strong, we can predict the flare class itself.
-However, this exacerbates the boundary issue, as it now occurs at multiple boundaries.
-Many works use image predictors to predict binary flare strength or flare classes.
+On the other hand, multi-channel images focused on active regions [Jin et al., 2026, in prep] must not only account for potential concurrent temporal effects (i.e., when active regions share timelines) but also for varying image resolutions, as active regions vary in size and shape.
+To address this, works such as Pandey et al.
+(2024) perform a rectangular crop that maximizes unsigned flux (a SHARP parameter).
+However, other work, such as Sun et al.
+(2023), employs a similar yet distinct approach by performing a crop centered on the polarity inversion line (PIL, Schrijver (2007)).
+PILs are typically the main source of flare activity, as shown in many works e.g., Zirin and Wang (1993); Green et al.
+(2018); Sun et al.
+(2021), so such crops can not only keep predictive power high but also allow for succinct comparisons of features (such as topological and spatial features) that the model detects with the derived PIL for potential interpretability results.
+Multi-channel images of active regions are commonly adopted in flare prediction.
+Typically, machine learning models are trained to forecast whether a flare will be “strong” within a set time period: a binary prediction with response being $\geq M$ (or $\geq C$) flares.
+Despite its simplicity, binary classification still presents challenges such as class imbalance and ill-defined boundaries between flare categories, particularly near the chosen threshold (e.g., just below or above the M-class).
+Alternatively, models can try to predict the exact flare class, but this introduces similar boundary problems across multiple class thresholds.
+Many studies use image predictors to predict binary flare classes.
 We briefly review a subset of these.
-Focusing on works that use only 2D magnetogram imaging data, Park2018 trains a deep Convolutional Neural Network (CNN) on full-disk SDO/HMI magnetograms for binary predictions of C+ flares.
-Meanwhile, huang2018deep and zheng2019solar train CNNs to do multi-class prediction of C-, M-, and X-class flares using full-disk magnetograms.
-On the other hand, li2022knowledge and sun2022predicting train CNNs on resized HMI magnetograms focused on active regions.
-Beyond using only 2D magnetogram data, jonas2018flare uses Gabor filters and other techniques on a subset of the AIA imaging channels to train a linear model to predict M+ binary flare outcomes.
-Additionally, nishizuka2018deep trains a deep neural network on AIA imaging channels to predict the binary responses for both C+ and M+ flares.
-For works combining both AIA and HMI images, nishizuka2021operational trains a CNN for binary predictions using both C+ and M+ decision rules on full-disk HMI and AIA images, and sun2023tensor proposes a tensor Gaussian process with contraction to combine AIA and HMI images to predict strong flares (M-class) versus weak flares (B-class).
-Image Predictors, Continuous Response
-Instead of predicting flare classes, another generalization for flare prediction is to directly predict the peak X-ray flux intensity of a flare, now treating the task as a regression problem rather than classification.
-There is not as much work done along this line as regression is inherently more difficult than classification, but we still provide a brief review here.
-In muranushi2015ufcorin, the authors use 2D line-of-sight full-disk magnetogram data and then perform a wavelet transform to train a time-series regression model to predict the peak flux intensity of flares.
-On the other hand, sun2023tensor uses both AIA and HMI imaging data aligned on active regions to perform a tensor regression to predict peak flux intensity.
-In these works, the models predict a continuous response, but the predictions are then converted into “classification” outputs using flare-class cutoffs in order to use commonly adopted classification metrics such as accuracy and skill scores.
-However, a work with which they do not bin their continuous predictions is in vandersande2025thes, which investigates key challenges and best practices that should be kept when treating flare prediction as strictly a regression problem.
-Vector Predictors, Binary Response
-In addition to high-resolution images that pose challenges to model training, many flare-forecasting studies rely on input data vectors that are derived from the summary statistics or data-driven features computed from the images.
-Space-Weather HMI Active Region Patches (SHARPs)(SHARPs), derived from photospheric magnetic field data taken by the HMI aboard the SDO, provide 25 magnetic and geometric summary parameters.
-SHARPs are widely used across flare, SEP, and CME forecasting applications.
-Similar to SHARPs, Space-Weather MDI Active Region Patches (SMARPs)(SMARPs) are derived from maps of the solar surface magnetic field taken by the Michelson Doppler Imager (MDI) aboard the Solar and Heliospheric Observatory (SOHO).
+Focusing on works that use only 2D magnetogram imaging data, Park et al.
+(2018) trains a deep Convolutional Neural Network (CNN) on full-disk SDO/HMI magnetograms for binary predictions of C+ flares.
+Meanwhile, Huang et al.
+(2018) and Zheng et al.
+(2019) train CNNs to perform multi-class predictions of C-, M-, and X-class flares using full-disk magnetograms.
+On the other hand, Li et al.
+(2022) and Sun et al.
+(2022) train CNNs on resized HMI magnetograms focused on active regions.
+Beyond using only 2D magnetogram data, Jonas et al.
+(2018) applies Gabor filters and other techniques to a subset of AIA imaging channels to train a linear model to predict $\geq M$ binary flare outcomes.
+Furthermore, Nishizuka et al.
+(2018) trains a deep neural network on AIA imaging channels to predict the binary responses for both $\geq C$ and $\geq M$ class flares.
+For works combining both the AIA and HMI images, Nishizuka et al.
+(2021) trains a CNN for binary predictions using both $\geq C$ and $\geq M$ decision rules on full-disk HMI and AIA images, and Sun et al.
+(2023) proposes a tensor Gaussian process regression with contraction operators to combine AIA and HMI images to predict strong flares ($\geq$ M-class) versus weak flares ($\leq$ B-class).
+1.1.2 Image Predictors and Continuous Responses
+Instead of predicting flare classes, another generalization of flare prediction is to directly predict the peak X-ray flux of a flare, treating the task as a regression problem rather than a classification problem.
+There is less work in this area because regression is inherently more difficult than classification, but we still provide a brief review here.
+In Muranushi et al.
+(2015), the authors use 2D line-of-sight full-disk magnetogram data and then perform a wavelet transform to train a time-series regression model to predict the peak flux intensity of flares.
+On the other hand, Sun et al.
+(2023) uses both AIA and HMI imaging data aligned on active regions to perform a tensor regression to predict peak flux intensity.
+In these works, the models predict a continuous response, but the predictions are then converted into “classification” outputs using flare-class cutoffs to enable the use of commonly adopted classification metrics such as accuracy and skill scores.
+However, a work in which they do not bin their continuous predictions is van der Sande and Muñoz-Jaramillo (2025), which investigates key challenges and best practices to consider when treating flare prediction as a strictly regression problem.
+1.1.3 Vector Predictors and Binary Responses
+In addition to high-resolution images that pose challenges for model training, many flare-forecasting studies rely on input data vectors derived from summary statistics or data-driven features computed from the images.
+Space-Weather HMI Active Region Patches (SHARPs)(Bobra et al., 2014), derived from photospheric magnetic field data taken by the HMI aboard the SDO, provide 25 magnetic and geometric summary parameters.
+SHARPs are widely used for flare forecasting and have recently demonstrated success in SEP predictions (Yu et al., 2025).
+Similar to SHARPs, Space-Weather MDI Active Region Patches (SMARPs)(Bobra et al., 2021) are derived from maps of the solar surface magnetic field taken by the Michelson Doppler Imager (MDI) aboard the Solar and Heliospheric Observatory (SOHO).
+Sun et al.
+(2022) adopts the SMARPs for flare forecasting.
 The binary classification of strong flares (such as $\geq$ M-class) against weak flares (such as $\leq C$ or $\leq B$ class) with multi-dimensional time series predictors is the most prevalent in the data-driven solar flare forecasting literature.
-Previous work on machine learning and deep learning models trained on multidimensional predictors extracted from photospheric magnetic field data (such as SHARPs) shows that carefully engineered vector physical summaries can already achieve competitive performance (bobra2015sola; florios2018forecasting; liu2019predicting; chen2019iden; nishizuka2018deep).
+Previous work on machine learning and deep learning models trained on multidimensional predictors extracted from photospheric magnetic field data (such as SHARPs) shows that carefully engineered vector physical summaries can already achieve competitive performance (Bobra and Couvidat, 2015; Florios et al., 2018; Liu et al., 2019b; Chen et al., 2019a; Nishizuka et al., 2018; Wang et al., 2020b; Nishizuka et al., 2021).
 Vector predictors are not limited to features derived from images.
-(verma2026optimal) studies the probability of a strong flare event as an extreme-event prediction using soft X-ray flux time-series vectors.
-Vector Predictors, Continuous Response
-Flares are classified based on their peak intensity, which is defined to be the peak value of the GOES X-ray flux during the flare; for example, M-class flares have a peak intensity that is at least ${10}^{-5}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$ and less than ${10}^{-4}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$, and X-class flares have a peak intensity that is at least ${10}^{-4}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$ (goes8-15_readme).
-A typical formulation of the flare forecasting problem is predicting whether an M+ flare will occur in a future 24-hour period (leka2018sola).
+(Verma et al., 2026) studies the probability of a strong flare event as an extreme-event prediction using soft X-ray flux time-series vectors.
+1.1.4 Vector Predictors and Continuous Responses
+Flares are classified based on their peak intensity, which is defined to be the peak value of the GOES X-ray flux during the flare.
+For example, M-class flares have a peak intensity that is at least ${10}^{-5}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$ and less than ${10}^{-4}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$, and X-class flares have a peak intensity that is at least ${10}^{-4}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$ (Mothersbaugh III et al., 2023).
+A typical formulation of the flare forecasting problem is predicting whether an M+ flare will occur in a future 24-hour period (Leka and Barnes, 2018a).
 In this formulation, the outcome is binary.
 Table 1 provides a justification for an alternative formulation in which the outcome is continuous.
-The table is based on the SWAN-SF benchmark dataset, which contains various kinds of data, including flare and GOES X-ray flux data, for the period 2010-2018 (angryk2020multivariate).
-For each date between 1 May 2010 and 29 December 2018, we checked whether an M+ flare occurred and whether the maximum X-ray flux was at or above the M-class threshold of ${10}^{-5}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$.
+The table is based on the Space Weather Analytics for Solar Flares (SWAN-SF) benchmark dataset, which contains various kinds of data, including flare and GOES X-ray flux data, for the period 2010-2018 (Angryk et al., 2020).
+For each date between May 1, 2010 and December 29, 2018, we checked whether an M+ flare occurred and whether the maximum X-ray flux was at or above the M-class threshold of ${10}^{-5}\text{\,}\mathrm{W}\,\mathrm{m}^{-2}$.
 All but four dates either had both an M+ flare and an M+ maximum X-ray flux or had neither.
 This shows that predicting whether an M+ flare will occur during a 24-hour window is essentially equivalent to predicting whether the maximum X-ray flux during the window will be above the M-class threshold.
-Predicting the latter event can be done by forecasting the maximum X-ray flux, which is a continuous outcome.
+Predicting the latter event can be achieved by forecasting the maximum X-ray flux, which is a continuous outcome.
+See Verma et al.
+(2026) for more discussions.
 Date Had M+ Flare?
@@
 No 4 (0.15%) 2,146 (83%)
-Table 1: The joint distribution of the M+ maximum flux and flare indicators for dates between 1 May 2010 and 29 December 2018, inclusive.
-The SWAN-SF benchmark dataset described in angryk2020multivariate was used to compute the counts.
+Table 1: The joint distribution of the M+ maximum flux and flare indicators for dates between May 1, 2010 and December 29, 2018, inclusive.
+The SWAN-SF benchmark dataset described in Angryk et al.
+(2020) was used to compute the counts.
 Forecasting a continuous outcome has several advantages over forecasting a binary outcome.
-First, a prediction of a binary outcome is a prediction of whether an event of interest will happen, but not of how severe the event will be if it does happen.
-A prediction of a continuous outcome is a prediction of both occurrence and severity.
-Second, the flare class thresholds are arbitrary; there is no substantive difference between the strongest flares in one class and the weakest flares in the next higher class.
-If ignored, the lack of a clear distinction between adjacent flare classes can make it difficult for a binary forecasting method to perform well, because the method may have trouble correctly predicting flares near a class boundary.
-Several works use a continuous outcome instead of a binary outcome.
-For example, vandersande2025thes models the maximum soft X-ray flux in a 24-hour window using linear regression, random forests, multilayer perceptrons (MLPs), and convolutional neural networks (CNNs).
-boucheron2015pred employs support vector regression (SVR) trained on magnetic features extracted from MDI images, using a linearly spaced continuous-valued label vector to represent flare size.
-jiao2020solar develops a mixed LSTM model to forecast the maximum flare intensity, and considers classification models built on top of the regression.
-In the sequel, we will describe issues that need to be addressed when using the various kinds of data mentioned previously and suggest ways to address those issues.
-In Section 2, we discuss response data, covering various data sources, quality, and a systematic comparison.
-In Section 3, we discuss predictor data, with image-based predictors being discussed in Section 3.1 and vector predictors being discussed in Section 3.2.
+First, a prediction of a binary outcome is a prediction of whether an event of interest will occur, but not of its severity if it does.
+A prediction of a continuous outcome involves predictions of both occurrence and severity (see e.g., Jiao et al., 2020b; Verma et al., 2026).
+Second, the flare class thresholds are arbitrary, meaning that there is no substantive difference between the strongest flares in one class and the weakest flares in the next higher class.
+If ignored, the lack of a clear distinction between adjacent flare classes can make it difficult for a binary forecasting method to perform well because the method may have trouble correctly predicting flares near a class boundary.
+Several studies use a continuous outcome rather than a binary outcome.
+For example, van der Sande and Muñoz-Jaramillo (2025) models the maximum soft X-ray flux in a 24-hour window using linear regression, random forests, multilayer perceptrons (MLPs), and convolutional neural networks (CNNs).
+Boucheron et al.
+(2015) employs support vector regression (SVR) trained on magnetic features extracted from MDI images, using a linearly spaced continuous-valued label vector to represent flare size.
+Jiao et al.
+(2020b) develops a mixed LSTM model to forecast the maximum flare intensity, and considers classification models built on top of the regression.
+1.2 Outline
+In the sequel, we will describe the issues that arise when using the various types of data mentioned above in Section 1.1 and suggest ways to address them.
+In Section 2, we focus on response data, covering various data sources, quality, and a systematic comparison.
+In Section 3, we focus on predictor data, with image-based predictors discussed in Section 3.1 and vector predictors discussed in Section 3.2.
 Finally, in Section 4, we show how predictive performance can be degraded if the issues we raise aren’t addressed.
+We conclude in Section 5 with take-home messages on our numerical comparison results, clarifying strengths and limitations of different data products.
 2 Flare Response: Data Source, Quality, and Processing
-2.1 Sources of Data Products for Flare Forecasting
-2.1.1 GOES X-ray Flare
+2.1 Sources of Data Products for Flare Events
+2.1.1 GOES X-ray Flare Data
+Geostationary Operational Environmental Satellites (GOES) are operated by NOAA (National Oceanic and Atmospheric Administration ), and observe the Sun in X-ray spectrum.
+GOES flare event data, 1-minute average X-Ray flux data, are one of the most commonly used resources in flare forecasting tasks and is available from multiple organizations.
+The GOES X-Ray Sensor (XRS) measurements have been a crucial component of space weather operations since 1975.
+XRS measurements are in two bandpass channels commonly referred to as the XRS-A (0.05-0.4 nm) and XRS-B (0.1-0.8 nm).
+A sudden yet persistent increase in flux is detected as a likely flare.
+For GOES-R (GOES 16-18) data, the detection algorithm uses a variety of criteria, including an exponential fit, to determine that a flare has begun.
+Specifically, the algorithm collects a small block of 1-minute averaged XRS-B flux data and smooths the data by a small boxcar filter.
+The algorithm then checks whether the flux shows signs of a significant rise by evaluating two key features: the presence of an inflection point, where the curve transitions to a faster upward slope, and whether the rise exceeds a defined threshold based on the standard deviation of earlier values.
+If these conditions are met, it fits the smoothed data to an exponential function.
+The exponential curve must also show a strong, concave upward shape based on the configurable parameters.
+If all these criteria are satisfied, the algorithm declares a flare start.
+The flare peak time is recorded when the flux reaches its maximum.
+The end time of a flare is declared when the X-ray flux has declined to halfway between the flare peak and the background level (Machol et al., ).
 […]
@@
 Not every local maximum in flux will be labeled as one flare event.
-The flux and the flare events are the Science-Quality data processed by NCEI.
-A notable change between the GOES-R (GOES 16-18) and previous GOES data is that the operational (near-real-time) data prior to GOES-16 had scaling factors applied by SWPC to adjust the GOES 8-15 irradiances to match fluxes from GOES-7.
-Thus, the flare index before GOES-16 was based on the operational irradiances with scaling factors.
-To get the true irradiances, the scaling factor of 0.7 (for the XRS-B channel) had to be removed (see Section 2.2).
-The operational flare data based on the operational 1-minute X-Ray flux is prepared and stored by SWPC.
-SWPC provides near-real-time and recently observed X-ray flux data, solar and geomagnetic indices and solar event reports created from preliminary reports.
-Specifically, the SWPC offers Solar Events Lists (SEL) from 1996 to the present, accessible via its public warehouse (see ftp://ftp.swpc.noaa.gov/pub/warehouse/).
-The SEL provides daily preliminary lists of solar events, such as bright surges on the limb, filament disappearances, optical flares observed in H-alpha, X-ray flares, and others.
-By selecting the XRA event type, users can access operational X-ray flare event data dating back to 1996 (SEL).
+The flux and the flare events are the Science-Quality data processed by the NOAA National Centers for Environmental Information (NCEI).
+A notable difference between the GOES-R (GOES 16-18) and previous GOES data is that the operational (near-real-time) data prior to GOES-16 had scaling factors applied by the SWPC (Space Weather Prediction Center) to adjust the GOES 8-15 irradiance to match fluxes from GOES-7.
+Thus, the flare index before GOES-16 was based on the operational irradiance with scaling factors.
+To obtain the true irradiance, the scaling factor of 0.7 (for the XRS-B channel) have to be removed (see Section 2.2).
+Operational flare data based on the 1-minute X-Ray flux is prepared and stored by the SWPC.
+SWPC provides near-real-time and recently observed X-ray flux data, solar and geomagnetic indices, and solar event reports created from preliminary reports.
+Specifically, SWPC offers Solar Events Lists (SEL) from 1996 to the present, accessible through its public warehouse (see ftp://ftp.swpc.noaa.gov/pub/warehouse/).
+The SEL provides daily preliminary lists of solar events, including bright surges on the limb, filament disappearances, H-alpha optical flares, X-ray flares, and others.
+By selecting the XRA event type, users can access operational X-ray flare event data dating back to 1996 (SWPC, ).
 We refer to the flare event lists from the SEL as the SWPC-FTP list.
@@
 Additionally, the warehouse contains the daily SRS (Solar Region Summary) report, which provides detailed information about the solar active regions, including their positions.
-Science-Quality XRS data are produced by NOAA’s National Center for Environmental Information (NCEI) and differ from the operational products used at SWPC in that the data have been preprocessed and incorporate retrospective fixes for issues and outages in the operational product.
+Science-Quality XRS data is produced by NOAA’s National Center for Environmental Information (NCEI) and differ from the operational products used at SWPC in that the data have been preprocessed and incorporate retrospective fixes for issues and outages in the operational product.
 Specifically, NCEI has reprocessed GOES 8-18 operational flare summary data to produce science-quality flare summary data (see Section 2.2).
 The GOES 1-7 data will be completed in 2026.
-The Science-Quality flare summary data is based on the 1-minute Science-Quality X-Ray irradiances data.
+The Science-Quality flare summary data are based on 1-minute Science-Quality X-ray irradiance data.
 We refer to this flare catalog as the Science-Quality list in this paper.
@@
 However, NCEI provides Science-Quality flare location information for GOES-R (GOES 16-18) data.
-The flare location on the solar disk is determined based on the measurements from the high flux XRS-B2 quad diode detector and was calibrated and validated with comparisons to flare locations from the Heliophysics Event Registry (goesr_l2_userguide).
-And this can be used to match to the ARs.
+The flare location on the solar disk is determined based on the measurements from the high flux XRS-B2 quad diode detector and was calibrated and validated with comparisons to flare locations from the Heliophysics Event Registry (https://www.lmsal.com/heksearch/, Machol et al.
+).
+And this can be used to match flares to the corresponding ARs.
+Table 2 summarizes the different GOES X-ray flare sources.
 Table 2: Overview of GOES X-ray Flare Sources.
@@
 SunPy-HEK 1975 – present GOES 1–16 / Yes Yes
-2.1.2 AIA Flares and SolarSoft Latest Events
-Besides GOES X-Ray flare, AIA flares are solar flares detected and characterized based on extreme ultraviolet (EUV) observations from the Atmospheric Imaging Assembly (AIA lemen2012atmospheric) aboard the Solar Dynamics Observatory (SDO pesnell2012solar).
-The detection process uses a peak-detection algorithm on binned AIA images, primarily in the 193 Åwavelength, to identify the start, peak, end time, and approximate location of the flare.
-Basically, the AIA-flares detection algorithm detects flares as intervals of positive derivative and negative derivative around a local maximum (Martens2012).
-SolarSoft Latest Events(Freeland2002) is another catalog that regularly detects and collects solar flare events.
-The catalog is primarily designed to assign spatial locations to GOES flare events by using the best available X-ray/EUV full-disk data.
+2.1.2 AIA Flare Catalog, SolarSoft Latest Events, and HEK Repository
+Besides GOES X-Ray flare, AIA flares are solar flares detected and characterized based on extreme ultraviolet (EUV) observations from the AIA (Lemen et al.
+(2012)) aboard SDO( Pesnell et al.
+(2012a)).
+The detection process uses a peak-detection algorithm applied to binned AIA images, primarily at 193 Å, to identify the start, peak, and end times and the approximate location of the flare.
+The AIA flare detection algorithm detects flares as intervals of positive derivative and negative derivative around a local maximum (Martens et al., 2012).
+SolarSoft Latest Events (Freeland, 2002) is another catalog that regularly detects and collects solar flare events.
+The catalog is primarily designed to assign spatial locations to GOES flare events using the best available full-disk X-ray/EUV data.
 Since 2010, AIA 94åhas been the primary data source, and currently, GOES-R series SUVI 93/94 Åserves as a backup when AIA data is unavailable.
-It is independent of specific instruments (personal communications with Dr.
+It is independent of specific instruments (based on personal communications with Dr.
 Samuel Freeland).
-2.1.3 HEK Repository
-The HEK is a comprehensive metadata system designed to catalog and provide access to solar events, facilitating efficient data retrieval for researchers (hek).
-HEK serves as an integrated event repository rather than a direct data source.
+The HEK (Heliophysics Events Knowledgebase) is a comprehensive metadata system designed to catalog and provide access to solar events, facilitating efficient data retrieval for researchers (Hurlburt et al., 2012).
+The HEK serves as an integrated event repository rather than a direct data source.
 It aggregates flare information from multiple sources.
-HEK incorporates AIA flare detections from automated feature-finding tools applied to 94 ÅEUV images.
+The HEK incorporates AIA flare detections from automated feature-finding tools applied to 94 ÅEUV images.
 It also ingests external sources, including the SWPC operational GOES X-Ray flare list and the SSW catalog.
-Additionally, the SunPy package allows users to access solar information through its HEK module.
-According to the official SunPy document for accessing the GOES flare list, as well as the open source code, users can retrieve the desired SWPC operational flare events from the HEK repository by setting the appropriate event type and observatory.
+The SunPy package enables users to access solar data via its HEK module.
+According to the official SunPy documentation and the open-source code, users can retrieve the desired SWPC operational flare events from the HEK repository by specifying the appropriate event type and observatory.
 However, the GOES flare list from the SunPy-HEK module exhibits several discrepancies compared to the flare list available in the SWPC open warehouse, particularly in event counts and Active Region (AR) number information.
 Between January 1, 1997, and July 21, 2024, 289 flares (1.14%) appear in the SunPy-HEK list but are absent from the SWPC-FTP list, while 817 flares (3.15%) recorded by SWPC are not found in the SunPy-HEK list.
-Additionally, 9842 flare events have differing AR numbers between the two lists.
+Moreover, 9842 flare events have differing AR numbers between the two lists.
 Figure 3 shows the cumulative flare intensity since January 1, 2010, for both the SunPy-HEK and SWPC-FTP flare lists.
 A notable disparity appears after 2022, due to 728 flare events recorded by SunPy-HEK that are missing, and more than 3,000 flares are mislabeled with AR number 0 in the SunPy-HEK list.
-The HEK operation team suggested that the difference in event counts may be due to the double retrieval of the same flare after SWPC revisions (personal communication).
+The HEK operation team suggested that the difference in event counts may be due to the double retrieval of the same flare after SWPC revisions (based on personal communications).
 Thus, we do not recommend using the SunPy-HEK GOES flare list, especially the GOES-R data, for research purposes.
@@
 The science-quality data have been reprocessed from the start of the mission to the present date and incorporate retrospective fixes for issues and outages in the operational product.
-The XRS L1b data contains high-cadence soft X-ray irradiance measurements, and the L2 data consists of higher-order products such as irradiance averages, flare event summaries, and flare location products.
+The XRS L1b data contain high-cadence soft X-ray irradiance measurements, and the L2 data consist of higher-order products such as irradiance averages, flare event summaries, and flare location products.
 Science-quality L2 data products are created from science-quality L1b data.
 Up to October 2025, science-quality flare summary products are available for GOES-R (GOES 16, 17, 18) and GOES 8-15 satellites, dating from 1995/01/03 to the present.
-1-minute flux average product, another commonly used L2 product, has been processed to science-quality for GOES 8-18 satellites.
-This paper focuses mainly on the flare summary product, which contains the most widely used flare response information in flare prediction tasks.
-The science-quality data XRS have been reprocessed with many corrections, including smoothness, accuracy, scaling factor consistency, bandpass corrections, and improvement of the data quality flag (goes8-15_readme).
+The 1-minute flux average product, another commonly used L2 product, has been processed to science-quality for GOES 8-18 satellites.
+This paper focuses primarily on the flare summary product, which contains the most widely used flare-response information for flare-prediction tasks.
+The science-quality data XRS have been reprocessed with numerous corrections, including smoothing, accuracy, scaling factor consistency, bandpass corrections, and improvements to the data quality flag (Mothersbaugh III et al., 2023).
 The SWPC scaling factor has the most crucial effect on the flare summary product.
-Science-quality XRS irradiances are provided in physical units of W/m2 (watts per square meter), while the operational data before GOES-16 had SWPC-adjusted scaling factors that adjusted the GOES 8–15 irradiances to match those of GOES 7.
+Science-quality XRS irradiances are provided in physical units of W/m2 (watts per square meter), while the operational data before GOES-16 had SWPC-adjusted scaling factors that adjusted GOES 8–15 irradiance to match those of GOES 7.
 Since the flare summary product is based on the L1b data using the flare detection algorithm, the threshold for each flare class increased by 42%, meaning that an individual solar flare before GOES-R needed to be 42% larger than a flare after GOES-R to be given the same level of magnitude.
 Thus, fewer flares were covered in the operational data set during GOES 8-15.
-The science-quality data were corrected so that they did not contain the SWPC scaling factors.
+The science-quality data were corrected to exclude the SWPC scaling factors.
 Figure 4 shows the flare magnitude inconsistency in the operational data.
@@
 Figure 4: The flare peak flux ratio of SWPC-FTP list to the NCEI science-quality list through 2020/01/01 - 2024/7/21.
-Before December 2019, the SWPC applied a rescaling factor such that the ratio is centered at 0.7, while the centered ratio is back to 1 as GOES-16 became the primary operational satellite.
+Before December 2019, the SWPC applied a rescaling factor so that the ratio was centered at 0.7, whereas the centered ratio returned to 1 as GOES-16 became the primary operational satellite.
 The colored rectangle shows the data availability time range for GOES 13-18, which differs from the serving time as the primary satellite providing operational data.
-GOES-16 became the primary geostationary satellite for its operations in December 2019.
+GOES-16 became the primary geostationary satellite in December 2019.
 Figure 5: Flare event numbers from NCEI Science-Quality data and SWPC-FTP data.
@@
 2.3 Cross-Comparison of Science-Quality, Operational, and SSW Flare Lists
-To further look at the differences of these flare lists, we then compared the Science-Quality, SWPC-FTP, and SSW flare lists.
+To further investigate the differences of these flare lists, we compared the Science-Quality, SWPC-FTP, and SSW flare lists.
 The Science-Quality catalog was obtained through the NCEI website.
 The SWPC-FTP catalog is directly downloaded from the SWPC open warehouse due to the inconsistency between the FTP archive and SunPy-HEK records.
-The SSW flares were downloaded from their web archive (Freeland2002).
-Additionally, we downloaded the Science-Quality flare location production from NCEI and augmented the Science-Quality list using the pinpoint location information.
-The period of interest covers from 2010-01-01 to 2024-12-31.
-NCEI provides the processed science-quality data for each instrument separately, and the data availability coverage of one instrument usually does not align with the time range when the instrument is in operational use.
+The SSW flares were downloaded from their web archive (Freeland, 2002).
+In addition, we downloaded the Science-Quality flare location dataset from NCEI and augmented the Science-Quality list with pinpoint location information.
+Finally, Science-Quality and SSW have pinpoint flare location coordinates, while SWPC-FTP and SSW have NOAA AR numbers.
+The period of interest is from 2010-01-01 to 2024-12-31.
+NCEI provides processed science-quality data for each instrument separately, and the data availability coverage of one instrument usually does not align with the time range during which the instrument is in operational use.
 For example, GOES-16 reached geostationary orbit on 2015-11-29 and started to provide flare event data since 2017-02-09, while GOES-16 did not become NOAA’s primary source of solar X-ray flux until 2019-12-09.
 GOES-16 was replaced by GOES-18 as the primary satellite in December 2024.
-During this time range, GOES-13, -14, -15, and -16 had become the primary satellites.
-We compiled the flare summary data from different instruments to get the Science-Quality flare list.
-Since the flare location product is only available for the GOES-R series and the flare location information is crucial in predictive model training, we include the whole time range of the GOES-16 since 2017-02-19.
-Before GOES-16, we compiled the science-quality data from different GOES instruments based on primary use status (goes1-15_readme).
+During this period, GOES-13, -14, -15, and -16 became the primary satellites.
+We compiled the flare summary data from different instruments to obtain the Science-Quality flare list.
+Since the flare location product is available only for the GOES-R series and flare location information is crucial for training predictive models, we include the full time range of GOES-16 since 2017-02-19.
+Before GOES-16, we compiled science-quality data from various GOES instruments based on their primary use status (Machol et al., ).
 We show the time range of the compiled science-quality flare data in Table 3.
@@
 In general, there have been two operational GOES satellites for each instrument, a primary and a secondary.
-Primary satellite data should be used over secondary satellite data whenever it is available.
+Primary satellite data should be used in preference to secondary satellite data whenever available.
 There will be gaps in good data for the primary satellites due to eclipses and in-flight calibrations.
-Satellites are listed in the table only when there are changes.
-Data courtesy: Kimberly Moreland
+Satellites are listed in the table only when changes occur.
+Data courtesy: Kimberly Moreland.
 Date Science-Quality Primary Secondary
@@
 2009-12-31 14 14 None
-During this time range, there are 33215 Science-Quality records, distributed into 139 X-flares and 2497 M-flares.
-We also downloaded 27586 SWPC-FTP and 33489 SSW flares.
-We apply minimum data cleaning to these lists, only ensuring the validity of start time, peak time, and flare magnitude features.
+During this time range, there are 33,215 Science-Quality records, distributed into 139 X-flares and 2,497 M-flares.
+We also downloaded 27,586 SWPC-FTP, and 33,489 SSW flares.
+We perform minimum data cleaning to these lists, only ensuring the validity of start time, peak time, and flare magnitude features.
 Table 4 presents the number of flares with different magnitudes in our compiled Science-Quality, SWPC-FTP, and SSW lists.
@@
 SSW 124 2194 19791 10023 1357 33,489
-For location information, Science-Quality and SSW have pinpoint flare location coordinates, and SWPC-FTP and SSW have NOAA AR numbers information.
 We then attempt to match each Science-Quality flare with corresponding events in the SWPC-FTP and SSW catalogs based on temporal characteristics (i.e., flare peak time) and flare intensity.
 Specifically, we define tolerance thresholds for both peak time and the logarithm (base 10) of the flare intensity.
-For a given flare, we first filter candidates from the comparison catalog whose absolute difference in peak time falls within the defined time tolerance window.
-We then further refine the matches by selecting only those flares whose difference in log10 intensity is within the specified threshold.
-In cases where multiple matches satisfy these criteria, the flare with the closest peak time is selected.
+For a given flare, we first filter the comparison catalog to exclude candidates whose absolute difference in peak time exceeds the defined time tolerance window.
+We further refine the matches by selecting only flares whose log10 intensity difference is within the specified threshold.
+When multiple matches satisfy these criteria, the flare with the closest peak time is selected.
 Figure 6 shows the fraction of matched flares under various tolerance settings.
@@
 The extent to which the SSW catalog inherits this bias is not explicitly documented.
-However, our analysis shows that the SSW catalog exhibits a level of similarity to SWPC-FTP that exceeds the similarity observed between any other pair of catalogs.
+However, our analysis shows that the SSW catalog exhibits greater similarity to SWPC-FTP than to any other pair of catalogs.
 To evaluate this, we first matched flares between the Science-Quality and SSW lists using only peak time with a strict tolerance of zero.
@@
 Before GOES-16 became operational, the intensity differences clustered around 0.15, consistent with the expected offset introduced by the SWPC scaling factor.
-After 2019-12-07, the differences are centered near zero, indicating the removal of this bias.
+After 2019-12-07, the differences are centered near zero, indicating that this bias has been removed.
 Together with the matching-proportion pattern shown in Figure 6, these results demonstrate that the SSW catalog is not free from the SWPC scaling prior to GOES-16 and, more generally, that SSW aligns more closely with SWPC-FTP than with the Science-Quality catalog.
-Therefore, the SSW list should be regarded and processed as an operational dataset.
+Therefore, the SSW list should be treated as an operational dataset.
 Consequently, for events occurring before 2019-12-09, we adjusted the flare log intensity for SWPC-FTP and SSW by removing the SWPC scaling factor in log scale.
-In this study, we adopt a peak time tolerance of 15 minutes and a log 10 intensity tolerance of 0.3 as the criteria for identifying corresponding flare events across different catalogs.
+In this study, we adopt a 15-minute peak-time tolerance and a 0.3 log10 intensity tolerance as criteria for identifying corresponding flare events across different catalogs.
 As shown in Figure 6, when the log intensity tolerance is set over 0.3, the fraction of matched flares remains relatively stable.
 Figure 8 shows the distribution of Science-Quality flares that fail to match an event in either the operational SWPC-FTP or SSW catalogs.
-Only 58 flares of class M or higher (2 X-flares) are unmatched, and the overall discrepancy decreases substantially after 2019-12-07, when GOES-R satellites became primarily operational.
+Only 58 flares of class M or higher (2 X-flares) are unmatched, and the overall discrepancy decreases substantially after 2019-12-07, when the GOES-R satellites became primarily operational.
 Figure 7: Distribution of $\log_{10}$-intensity differences between matched SWPC-FTP and SSW flares.
@@
 However, the NCEI Science-Quality flare summary product does not include NOAA AR numbers.
-It only provides precise flare locations for the GOES-R series (GOES-16/17/18/19) beginning on 2017-02-09, expressed in four coordinate systems: Stonyhurst/heliographic (longitude, latitude), Carrington coordinates, heliocentric radial coordinates, and helioprojective Cartesian (HPC) $(x,y)$ goes8-15_readme.
-For Science-Quality flares that successfully match to a SWPC-FTP or SSW flare, we directly adopt the AR number from the matched operational report.
+It only provides precise flare locations for the GOES-R series (GOES-16/17/18/19) beginning on 2017-02-09, expressed in four coordinate systems: Stonyhurst/heliographic (longitude, latitude), Carrington coordinates, heliocentric radial coordinates, and helioprojective Cartesian (HPC) $(x,y)$ Mothersbaugh III et al.
+(2023).
+For Science-Quality flares that are matched to a SWPC-FTP or SSW flare, we directly adopt the AR number from the matched operational report.
 However, approximately 37% of science-quality flares cannot be matched, or are matched to entries associated with an invalid AR number (0).
-To extend AR coverage for these cases, we use the pinpoint flare locations provided in the GOES-R series together with NOAA Solar Region Summary (SRS) reports.
+To extend AR coverage for these cases, we use the pinpoint flare locations provided by the GOES-R series, together with NOAA Solar Region Summary (SRS) reports.
 The SWPC SRS reports, available daily through the public FTP archive, provide detailed information for each numbered active region, including its position in Stonyhurst/heliographic coordinates.
 For every Science-Quality flare with valid HPC coordinates, we assign an AR number by identifying the nearest active region in the SRS report using the Euclidean distance in HPC coordinates:
-$\text{Distance}\;=\;(X_{\text{AR}}-X_{\text{Flare}})^{2}+(Y_{\text{AR}}-Y_{\text{Flare}})^{2}.$
-Among the flares occurring during the interval for which precise location information is available, 93% (11,783 events) can be assigned a nearest AR.
-For those flares that also have a nonzero AR number reported in SWPC-FTP, 70% share the same AR number, and more than 95% differ by $\leq 6$ units.
-Table LABEL:tab:nearest_ar_distance summarizes the Euclidean distance (in arcseconds) between flare and nearest AR.
-To ensure reliability, we retain only AR assignments with distance <250<250 arcsec, increasing the number of valid AR labels for approximately 9% of all flares in this period.
+$\text{Distance}\;=\;\sqrt{(X_{\text{AR}}-X_{\text{Flare}})^{2}+(Y_{\text{AR}}-Y_{\text{Flare}})^{2}}.$
+Among the flares occurring during the interval for which precise location information is available, 93% (11,783 events) can be assigned to a nearest AR.
+For those flares that also have a nonzero AR number reported in SWPC-FTP, 70% of the assigned flares share the same AR number with that of the SWPC-FTP, and more than 95% differ by $\leq 6$ units of distance in HPC coordinates.
+Table 5 summarizes the Euclidean distance (in arcseconds) between a flare and the nearest AR.
+To ensure reliability, we retain only AR assignments with a distance of <250<250 arcsec.
+We then assign the nearest AR number to flares that are either unmatched in SWPC-FTP or reported with AR number 0.
+As a result, 87% of GOES-R flares have valid AR labels, compared with only 77% before augmentation, representing an increase of approximately 10% of all flares in this period.
 For flares occurring before the availability of GOES-R pinpoint locations, AR numbers are assigned only when a valid match exists in either SWPC-FTP or SSW.
-Eventually, we assign valid AR numbers to 72% of all Science-Quality flares.
+Eventually, we were able to assign valid AR numbers to 72% of all Science-Quality flares.
 This augmented science-quality flare list is used in the model-training comparisons presented in Section 4.
-A complete data-processing pipeline is provided in the Appendix.
+A complete data-processing pipeline is provided in Appendix C.
 Table 5: Summary statistics for Euclidean distance (in arcsec) between Science-Quality flare and the nearest Active Region in the helioprojective Cartesian coordinate.
@@
 3.1 Imaging Predictors: HMI/SDO, AIA/SDO, and SOHO/MDI
-The imaging dataset of interest is the imaging data provided by the Atmospheric Imaging Assembly (AIA) (lemen2012atmospheric) and Helioseismic and Magnetic Imager (HMI) (scherrer2012helioseismic) instruments from the Solar Dynamics Observatory (SDO) since its launch in 2010 (pesnell2012solar).
+The imaging dataset of interest is the imaging data provided by the Atmospheric Imaging Assembly (AIA) (Lemen et al., 2012) and Helioseismic and Magnetic Imager (HMI) (Scherrer et al., 2012) instruments from the Solar Dynamics Observatory (SDO) since its launch in 2010 (Pesnell et al., 2012a).
 These instruments capture the full-disk images of the Sun in high spatial resolution (4096 $\times 4096$) with high temporal frequency as well.
 However, the data from HMI and AIA do not align up one-to-one as the AIA images have a larger field of view, resulting in different pixel sizes.
-There has been much work in processing the SDO instrument data to help facilitate its use in machine learning research due to these kinds of discrepancies.
-For example, galvez2019machine curates a data set by removing poor quality images, rescaling the images to a lower and more manageable resolution (512 $\times 512$), and applied corrections due to machine degradation and exposure issues.
-In particular, the data product that we focus on is tailored specifically for machine-learning based flare prediction research [Jin et al., 2024 in prep], which processes the existing HMI HARP data with the AIA images (in the 8 channels of 94, 131, 171, 193, 211, 304, 35, and 1600 Ångstroms) to make it the same field of view and coordinates.
-This data set will contain all B, C, M, and X-class flares from 2010-2024, starting 24 hours before the flare peak-time at a temporal cadence of 12 minutes.
-The flare peak time is described by the Geostationary Operational Environmental Satellite (GOES) which is operated by the United State’s National Oceanic and Atmospheric Administration (NOAA).
-The HMI instrument is used to study oscillations and the magnetic field of the photosphere of the Sun’s surface.
-The image contain 2D photospheric maps of the 3 orthogonal magnetic components.
-This is in contrast to the MDI instrument from SOHO, which only recorded 2D Line-of-sight imaging data for the magnetic field, which has been popular for previous machine-learning models.
-In particular, we are interested in the radial field component of the magnetic image $B_{r}$ channel of the HMI images.
-This orthogonal component of the magnetic field is not only where some of the summary statistics (like SHARPS that we will discuss further in Section 3.2) for the Sun is derived from, but also the magnetic polarity inversion lines (PIL) can be identified from .
-This is important as the PIL is known to be one of the most important features in triggering flares and other solar weather activity (sun2023tensor).
-One issue regarding the solar imaging data from the HMI and AIA that needs to be worried about is with regards to “near-limb” images that are taken.
-By “near-limb”, we typically mean the images that are taken of active regions that are on the edge, which is a potential data quality issue.
-As an example, refer to Table 6, which show the mean proportion of missingness for images near the limb and in the Disk center (which we define to have a longitude of above 60 degrees or less than -60 degrees).
-As we can see, there is a significantly higher proportion of missings when the images are near the limb, with a lot more variability.
-Thankfully, we have more images on the disk center that are of “higher-quality”,
+There has been extensive work on processing SDO instrument data to facilitate its use in machine learning research, given these discrepancies.
+For example, Galvez et al.
+(2019) curates a data set by removing poor quality images, rescaling the images to a lower and more manageable resolution (512 $\times 512$), and applying corrections due to machine degradation and exposure issues.
+In particular, [Jin et al., 2026 in prep] processes the existing HMI HARP data with the AIA images (in the 8 channels of 94, 131, 171, 193, 211, 304, 35, and 1600 Ångstroms) to make it the same field of view and coordinates, which is tailored specifically for machine-learning-based flare prediction research.
+This data set contains all B, C, M, and X-class flares from 2010-2024, starting 24 hours before the flare peak-time at a temporal cadence of 12 minutes.
+The flare peak time is described by the Geostationary Operational Environmental Satellite (GOES), which is operated by the United States’ National Oceanic and Atmospheric Administration (NOAA).
+The HMI instrument is used to study oscillations and the magnetic field of the Sun’s photosphere.
+The image contains 2D photospheric maps of the 3 orthogonal magnetic components.
+This contrasts with the MDI instrument on SOHO, which recorded only 2D line-of-sight magnetic-field images; HMI contains the radial field component of the magnetic image, $B_{r}$.
+This orthogonal component of the magnetic field is not only where some of the summary statistics (like the SHARPS that we will discuss further in Section 3.2) for the Sun are derived from, but also where the magnetic polarity inversion lines (PIL) can be identified from.
+This is important because the PIL is known to be a key driver of flares and other solar weather phenomena (Sun et al., 2023).
+One issue with the solar imaging data from the HMI and AIA concerns “near-limb” images, those images of active regions at the limb, which can introduce potential data-quality issues.
+Table 6 shows the mean proportion of missingness for images near the limb and in the Disk center (which we define to have a longitude of above 60 degrees or less than -60 degrees).
+As shown, the proportion of missing data is significantly higher when the images are near the limb, with greater variability.
+Figure 9 is an example near-limb image of AR11079.
+There are clear quality issues with this image: not only is it blurry and skewed, but the specific details and contrast are not very distinct.
+In contrast, Figure 10 from AR11121 shows a nearly perfectly centered AR.
+The quality difference is substantial, especially given that the specific details are readily discernible.
+As such, when dealing with these solar flare images, it is best to focus on images that are not on the limb for quality control.
+After such a procedure, we can curate a list of images for use in modeling or other data tasks.
 Region Mean Prop Missing Std. Dev. Total Image Counts
@@
 Table 6: Mean proportion of missingness for images on the limb versus disk center along with total image counts
-For a qualitative perspective of the difference between near limb images and not, Figure 9 is an example of an image of AR11079 of a near limb image.
-There are clear quality issues with this image, as not only is it blurry and askew, but the specific details and contrast are not very distinct.
-As such, there is limited information that can possibly be learned by such images.
-Figure 9: AR11079 Example Image at 2010-06-13T023100 which is on the limb with a max longitude that is around 80 degrees. It is off-center, blurry, and difficult to make out the specific details.
-On the other hand, we have Figure 10 which is of AR11121 when it is near perfectly centered.
-The quality difference is quite immense, especially as the specific details are quite easy to even visually decipher.
+Figure 9: AR11079 Example Image at 2010-06-13T023100 which is on the limb with a max longitude that is around 80 degrees. It is off-center, blurry, and difficult to discern specific details.
 Figure 10: AR11121 Example Image at 2010-11-11 04:36:00 which is a clear, centered image not on the limbs of the instrument.
-As such, when dealing with these solar flare images, for quality control is best to only deal with the images that are not on the limb.
-After such a procedure, we can then curate a list of images for the sake of any modeling or other data tasks.
-Another concern with the solar flare imaging data is the sporadic missingness we observe in some videos.
-For many analyses, we would like a continuous 24-hour sequence of an active region’s evolution prior to flare onset.
-In practice, such complete coverage is uncommon, because the active region must be identified early enough and remain in view for the entire 24 hours.
-In addition, rare AIA data gaps can occur which mean that even for regions currently being observed, some time intervals are not recorded, creating “gaps” in otherwise continuous video data.
-For a visualization to depict the potential negative effects of this data gap, for the task of X hour again prediction of extreme space weather events like solar flares or SEPs, it may be of interest to use historical imaging data of active regions of 24 hours (or 6 hours or anything of interest) before an eruption as training data for a model.
-However, if we use a flare list listing all flares from 2010-2024 and their associated active region and look at their available images starting from 1 hour away from peak time to 24 hours away, we can see that there is generally less data available the further away we are trying to forecast.
+Another concern with the solar flare imaging data is the sporadic missingness in some ARs.
+Ideally, a continuous 24-hour sequence of an active region’s evolution prior to flare onset would be used as the predictors for machine learning models.
+In practice, such complete coverage is uncommon because the active region must be identified early enough and remain in view for the entire 24 hours.
+In addition, AIA data gaps can occur, meaning that even for regions currently being observed, some time intervals are not recorded, creating gaps in otherwise continuous video data.
+Due to the cyclic nature of the Sun’s rotation as well.
+We will have gaps in observations of active region patches on the Sun because, when they rotate out of our view, we will no longer see them until they rotate back.
+This means that if there were any solar activity, such as the formation of a new active region or the end of one, the specific time and moment may not be captured.
+For the prediction of solar flares, it is desirable to use historical imaging data of active regions 24 hours before an eruption as training data.
+However, if we use a flare list listing all flares from 2010-2024 and their associated active regions, and examine images available from 24 hours before peak time to peak time, we find that there is generally less data available the further we attempt to forecast.
 This visualization is in Figure 11.
 Figure 11: Temporal coverage of on-disk flare images.
-(Left) Count of unique flares with available data not on the limbs as a function of nearest hour until flare onset.
+(Left) Count of unique flares with available data not on the limbs as a function of the nearest hour until flare onset.
 (Right) Stacked annual counts of unique flares that possess complete (non-limb images) 24-hour data coverage, classified by X-ray intensity (M/X, C, and B classes).
 3.2 Vector Predictors: SHARPs
-If a machine learning model uses images to predict, then it needs to be trained on a large volume of images in order to predict well.
+If a machine learning model uses images for prediction, it must be trained on a large volume of images to perform well.
 As stated in Section 3.1, HMI images are high-resolution.
-High-resolution images make for big files and are thus challenging to store in large numbers and to efficiently process in training pipelines.
+High-resolution images generate large files and are thus challenging to store in large quantities and to process efficiently within training pipelines.
 An alternative to using images is to use vectors of summary statistics computed from images; the SHARP dataset is based on this idea.
-The following description of the dataset draws primarily on (SHARPs).
+The following description of the dataset draws primarily on (Bobra et al., 2014).
 SHARP stands for Space-weather HMI Active Region Patch.
-An HMI Active Region Patch, or HARP, is a magnetic structure akin to an active region (HARPs), though a HARP can contain multiple NOAA active regions (SHARPs).
-A software pipeline automatically detects HARPs in HMI images and tracks them as they traverse the solar disk (HARPs).
+An HMI Active Region Patch, or HARP, is a magnetic structure akin to an active region (JSOC, ), though a HARP can contain multiple NOAA active regions (Bobra et al., 2014).
+A software pipeline automatically detects HARPs in HMI images and tracks them as they traverse the solar disk (JSOC, ).
 Every twelve minutes, the pixels within a HARP are used to calculate the vector of sixteen SHARP parameters, each of which describes some aspect of the vector magnetic field of the active region within the HARP.
-See Table 3 in SHARPs for the definitions and explanations of the SHARPs.
+See Table 3 in Bobra et al.
+(2014) for the definitions and explanations of the SHARPs.
 For example, the parameter USFLUX measures the total unsigned flux, and the parameter MEANJZD measures the vertical current density.
-Previous studies have demonstrated that the SHARP parameters have a fair degree of predictive power for forecasting flares (e.g., bobra2015sola; chen2019iden), validating the idea that summary statistics can be an effective alternative to images.
+Previous studies have demonstrated that the SHARP parameters exhibit fair predictive power for flare forecasting (e.g., Bobra and Couvidat, 2015; Chen et al., 2019a), supporting the notion that summary statistics can be an effective alternative to images.
 There are two types of HARP products: definitive (DEF-SHARPs) and near-real-time (NRT-SHARPs).
@@
 It is important to note that HARP identifiers are not consistent between the definitive and NRT products.
-The DEF-SHARPs dataset spans an additional two years beyond the availability of NRT-SHARPs (starting on 2010-05-01 versus 2012-09-14), resulting in approximately 2.3 million additional measurement samples that are absent in the NRT set.
-Even within the overlapping time period, definitive SHARPs include 912 more unique NOAA active regions (ARs) than the NRT product, with only 1043 ARs present in both datasets.
-Unless otherwise specified, the analyses that follow are based on the definitive SHARPs dataset.
+The DEF-SHARPs dataset spans two additional years beyond the availability of NRT-SHARPs (starting on 2010-05-01 rather than 2012-09-14), yielding approximately 2.3 million additional measurement samples not present in the NRT set.
+Even within the overlapping period, definitive SHARPs include 912 more unique NOAA active regions (ARs) than the NRT product, with only 1043 ARs present in both datasets.
 We compare these two products in terms of prediction model performance in Section 4.3.
-Many other quantities are calculated with the SHARP parameters, such as various longitudes and latitudes.
+Many other quantities are computed alongside the SHARP parameters, including various longitudes and latitudes.
 The bits of the QUALITY keyword indicate whether various issues, like eclipses, make the calculated values unreliable; see http://jsoc.stanford.edu/jsocwiki/Lev1qualBits for all the issues that can be flagged.
-Even if the QUALITY keyword does not indicate any problems, some SHARP parameters can still be missing values; at other times, the entire record, i.e., the SHARP parameters, longitudes and latitudes, QUALITY, etc., may be missing.
-Keywords that can only be computed when a magnetic field is present can have missing values during “padding intervals”; keywords can also have missing values for HARPs that are faint (HARPs).
-We do not know if these are the only possible causes of missingness.
+Even if the QUALITY keyword does not indicate any problems, some SHARP parameters can still be missing; at other times, the entire record, i.e., the SHARP parameters, longitudes and latitudes, QUALITY, etc., may be missing.
+Keywords that can only be computed when a magnetic field is present can have missing values during “padding intervals”.
+Keywords can also have missing values for HARPs that are faint (JSOC, ).
+We do not even know if these are the only possible causes of missingness.
 We investigated quality issues in the HARP and SHARP data for the period 1 May 2010 to 17 August 2023.
-The former date is the earliest date for which SHARP data is available (SHARPs).
+The former date is the earliest date for which SHARP data is available (Bobra et al., 2014).
 As stated earlier, for any HARP, a new record is created every twelve minutes.
@@
 We found that 526,817 (90%) of the times were within the lifetime of some HARP, and 56,024 (10%) were not within the lifetime of any HARP.
-The sequence of times can be partitioned into runs such that within each run, either every time is in some HARP lifetime or no time is in some HARP lifetime.
+The sequence of times can be partitioned into runs such that, within each run, either every time is in some HARP lifetime or no time is in some HARP lifetime.
 For each run, we computed its start and end times and its duration.
@@
 The first coordinate of a dot is the start time of the corresponding run, while the second coordinate is the run’s duration in days.
-Orange runs should not be used as there is no HARP data for them and no way to impute it.
+Orange runs should not be used, as there is no HARP data for them and no way to impute it.
 It seems reasonable to train a model with at least a few months’ worth of data; there are only two blue runs that could be divided into training sets with that much data.
@@
 A HARP record can have one or more problems.
-A record can be missing values for all SHARP parameters; we call such records missing.
-A record can be missing values for some, but not all SHARP parameters; we call such records incomplete.
+A record can have missing values for all SHARP parameters and we call such records missing.
+A record can be missing values for some, but not all SHARP parameters and we call such records incomplete.
 We flagged records with these problems.
@@
 Another problem is that for certain records, some SHARP parameter is infinite.
-Yet another is that some records are marked as low-quality.
+Another is that some records are marked as low quality.
 We flagged records with these problems.
@@
 Image quality degrades near the limb of the Sun, so the problems for a record could be due to the HARP being far from the central meridian.
-As Table 8 shows, most records don’t have any of the problems described above, but a sizable fraction have some problem.
+As Table 8 shows, most records do not have any of the problems described above, but a sizable fraction have some problem.
 Only 0.1% of records have more than one problem.
-The most important subsets of problematic records consist of records that are only missing (8.1% of records), records that are only low-quality (5.4% of records), and records that are only incomplete (1.3% of records).
+The most important subsets of problematic records are those that are only missing (8.1% of records), only low-quality (5.4% of records), and only incomplete (1.3% of records).
 Records with an infinite value are virtually nonexistent.
@@
 85% of records have no problems, and less than 1% have multiple problems.
-Of those records that do have problems, most are either missing or of low-quality.
+Of those records that do have problems, most are either missing or of low quality.
 4 Comparison of Data Products from a Flare Prediction Perspective
@@
 Note that we do not exhaust the available machine learning models and do not aim to optimize reported skill scores here, since, as discussed in the Introduction, (i) when properly trained, commonly used machine learning models do not generate significantly different results, and (ii) the skill scores of machine learning models depend heavily on sample selection and sample split schemes.
-Therefore, we use common practice for the choice of machine learning model and sample split here to demonstrate the differences in the various data products’ effects on the trained model’s performance.
+Therefore, we follow standard practice in selecting the machine learning model and sample split to demonstrate the effects of the various data products on the trained model’s performance.
 4.1 Data Preparation and Model Setting
-To assess how data quality affects flare forecasting performance, we consider two sources of flare labels (the SWPC-FTP list and the Science-Quality list) and two versions of predictor data (HMI near-real-time SHARPs and definitive calibrated SHARPs), see Table 9.
+To assess how data quality affects flare forecasting performance, we consider two sources of flare labels (the SWPC-FTP list and the Science-Quality list) and two versions of predictor data (HMI near-real-time SHARPs and definitive calibrated SHARPs), as shown in Table 9.
 We evaluate two representative model families: a deep learning architecture (LSTM) and a conventional statistical model (logistic regression with PCA, i.e., principal component analysis).
@@
 Table 9: Combinations of flare response datasets and SHARP predictor data used in model comparison. For every combination, we train an LSTM-based classifier and a logistic regression model.
-We follow the data preparation and sample construction methods described in  (Jiao2020lstm).
+We follow the data preparation and sample construction methods described in  (Jiao et al., 2020a).
 The input data consists of 24-hour time series extracted from the full time series of SHARP parameters of ARs.
@@
 Consequently, the target labels can be categorized as ”A,” ”B,” ”C,” ”M,” and ”X”.
-In this paper, we focus on a binary classification problem that distinguishes strong(M/X) from Weak (A/B) flares.
+We focus on a binary classification problem that distinguishes strong(M/X) from Weak (A/B) flares.
 We exclude the C-class category because its physical significance is ambiguous.
-C-class flares may be comparable to either small M-class flares or strong B-class flares, making it an unreliable separator boundary and thus vague the training process (chen2019solar).
+C-class flares may be comparable to either small M-class flares or strong B-class flares, making it an unreliable separator boundary and thus vague the training process (Chen et al., 2019b).
 In this study, we adopt two models: an LSTM model with a sigmoid output layer and a logistic regression model.
-The LSTM-based model is widely used in solar flare forecasting for handling time-series inputs ((chen2019iden), (Liu2019lstm), (Wang2020lstm)) and has been shown to outperform other machine learning models ((Wang2020lstm)).
+The LSTM-based model is widely used in solar flare forecasting for handling time-series inputs ((Chen et al., 2019a), (Liu et al., 2019a), (Wang et al., 2020b)) and has been shown to outperform other machine learning models ((Wang et al., 2020b)).
 The network consists of a two-layer LSTM with a hidden size of 30 units and dropout of 0.3 between layers, followed by a ReLU activation and a fully connected layer that maps the final (or truncated) hidden representation to a single scalar logit.
 This logit value is interpreted as the pre-sigmoid score for the probability of a strong flare.
-We also include logistic regression due to its simplicity, which in fact results in stability for trained models, and interpretability.
+We include logistic regression due to its simplicity, which in fact results in stability for trained models and interpretability.
 Given the $k$-hour forecasting window used to extract SHARP parameters, the input of the LSTM model for each target is a $5k\times 20$ matrix, where each row represents a 20-dimensional vector of SHARP parameters.
@@
     For the logistic regression model, we then apply principal component analysis (PCA) to reduce dimensionality, retaining 98% variance.
-For the 24-hr forecasting window, 20 principal components already cover over 98% variance.
+For the 24-hr forecasting window, 20 principal components already account for over 98% of the variance.
 The projection coefficients for these principal components are derived from the training set.
   3. 3.
-    The LSTM model is optimized using the Adam optimizer with a binary cross-entropy loss.
+    The LSTM model is optimized using the Adam’s optimizer with a binary cross-entropy loss.
   4. 4.
-    The logistic regression model is fitted with $L_{2}$ regularization. The penalty strength is optimized with 5-fold cross-validation and the ROC-AUC scoring objective.
+    The logistic regression model is fitted with $L_{2}$ regularization.
+The penalty strength is optimized with 5-fold cross-validation and the ROC-AUC (Receiver Operating Characteristic and Area Under the Curve) scoring objective.
 After each bootstrap model is trained, we evaluate probabilistic predictions on the held-out validation set and determine the optimal decision threshold that maximizes the True Skill Statistic (TSS) over a grid of candidate thresholds in $(0,1)$.
 The optimal threshold and corresponding validation TSS scores are recorded for every bootstrap sample.
-To properly evaluate the performance of the machine learning algorithms, we split the samples (time sequences of SHARP parameters and their corresponding maximum flare magnitudes) into training and testing sets.
+To evaluate the performance of the machine learning algorithms, we split the samples (time series of SHARP parameters and their corresponding maximum flare magnitudes) into training and test sets.
 Note that splitting by Active Region (AR) or HARP is crucial to prevent information leakage from the training set to the testing set.
 In this study, all training/testing splits are performed chronologically by HARP, which mimics the operational setting but may not result in the highest skill scores.
-See wang2020solar for detailed discussions on sample splitting schemes and their impacts on flare prediction results.
-It is important to note that varying definitions of positive and negative samples make direct comparisons of results from existing literature challenging, as discussed in chen2024solar.
+See Wang et al.
+(2020a) for detailed discussions on sample splitting schemes and their impacts on flare prediction results.
+It is important to note that varying definitions of positive and negative samples make direct comparisons of results from existing literature challenging, as discussed in Chen et al.
+(2024).
 The purpose of this paper is not to propose a universal method for sample construction but rather to compare model performance across different datasets.
-We also demonstrate numerically that the differences arising from data inconsistency or quality are not affected by the choice of the sample construction method.
+We also demonstrate numerically that differences arising from data inconsistencies or quality are unaffected by the choice of sample construction method.
 4.2 Science-Quality versus Operational “Response”
-To quantify the effect of flare-label quality on the prediction models’ skill scores, we compare the LSTM model and the logistic regression model for 6/12/24 forecasting trained using Science-Quality and SWPC-FTP flare list while fixing the SHARPs parameter inputs at definitive SHARPs.
-Box plots of six commonly adopted skill scores (TSS, HSS, POD, F1, FAR, ACC, see, e.g., chen2019iden for definitions) for the 24-hour forecasting windows are shown in Figure 13 and Figure 14 for the LSTM model and the logistic regression model, respectively.
+To quantify the effect of flare-label quality on the skill scores of the prediction models, we compare an LSTM model and a logistic regression model for 6/12/24 forecasting, trained using Science-Quality and SWPC-FTP flare lists, while using the definitive SHARPs parameter as inputs.
+Box plots of six commonly adopted skill scores (TSS, HSS, POD, F1, FAR, ACC, see, e.g., Chen et al.
+(2019a) for definitions) for the 24-hour forecasting windows are shown in Figure 13 and Figure 14 for the LSTM model and the logistic regression model, respectively.
 The 6-hour and 12-hour window prediction results exhibit consistent trends as these.
-The numerical results for all different model setups are given in Tables in Appendix B.
-We trained models using data from Solar Cycle 24 (5/1/2010–12/31/2019) and tested them with different time intervals within Solar Cycle 25.
+The numerical results for all model setups are presented in the Tables in Appendix B.
+We trained models using data from Solar Cycle 24 (5/1/2010–12/31/2019) and tested them using different time intervals in Solar Cycle 25.
 The evaluation spans four distinct testing intervals that correspond to different phases of Solar Cycle 25.
@@
 The 2023–2024 interval corresponds to the solar maximum, featuring intense flare activity and a high positive sample ratio (often exceeding 0.8).
-Finally, the combined 2020–2025 evaluation window averages performance across the full range of solar activity and reflects the overall performance of the model with a balanced test set.
+Finally, the combined 2020–2025 evaluation window averages performance across the full range of solar activity and reflects the model’s overall performance with a balanced test set.
 Across the test periods when the solar activity is relatively low, the models trained with SWPC-FTP operational flare list achieve higher predictive skill scores than those trained with Science-Quality labels.
@@
 During the solar minimum (20200101-20211231), the logistic regression model improves from a TSS of 0.15 [0.11,0.18] (Science-Quality) to 0.37 [0.35,0.38] (SWPC-FTP), and the LSTM model improves from 0.42 [0.16,0.59] to 0.57 [0.40,0.69].
-Similar margins in performance win appear in other metrics.
+Similar margins in performance wins are observed in other metrics.
 However, for the LSTM model, the difference is not substantial.
@@
 During the high-activity period (2023–2024) when the positive-sample ratio exceeds 0.85, the Science-Quality-trained models achieve better performance except for the TSS score of the logistic models.
-Across all the test periods, models trained with SWPC-FTP list get lower FAR scores, demonstrating that models trained with operational flare lists are better at identifying flares without increasing false alarms.
+Across all test periods, models trained with SWPC-FTP lists achieve lower FAR, demonstrating that models trained with operational flare lists are better at identifying flares without increasing false alarms.
 Figure 13: Box plots of six skill scores for different testing year choices (given in the x-axis) in the solar cycle 25 of the M/X versus A/B classification task by the LSTM model trained on definitive SHARPs as inputs with a 24-hour forecasting window.
 The blue boxes represent the Science-Quality flare list, while the orange boxes represent the SWPC-FP operational flare list.
-The dashed lines indicate the positive sample proportions in the test set, corresponding to the Y axis on the right-hand-side of each sub-figure.
+The dashed lines indicate the positive sample proportions in the test set, corresponding to the Y axis on the right-hand side of each sub-figure.
 Note that a smaller FAR means better performance.
@@
 The upper and lower error bars are at $Q_{3}+1.5(Q_{3}-Q_{1})$ and $Q_{3}-1.5(Q_{3}-Q_{1})$, respectively.
-The small dots outside of the boxes show the data points outside the error bars (outliers).
-The mean value and median are calculated including the outliers.
+The small dots outside the boxes indicate data points that fall outside the error bars (outliers).
+The mean value and median are calculated, including the outliers.
 Figure 14: Same as Figure 13 except the model being the logistic regression.
@@
 In addition, the performance differences between LSTM and logistic models are noticeably larger during solar maximum than during solar minimum.
-This suggests that the more complex flare environment during solar maximum makes model choice more critical and that simpler models as logistic regression can have an advantage.
+This suggests that the more complex flare environment during solar maximum makes model choice more critical and that simpler models, such as logistic regression, can have an advantage.
 (a) TSS, Solar Minimum
@@
 Figure 16 shows the yearly positive rate and positive-sample counts for both DEF-SHARPs and NRT-SHARPs.
-Despite having fewer valid observations, the NRT-SHARPs product exhibits a positive-rate pattern that closely matches the definitive product across all training years.
+Despite fewer valid observations, the NRT-SHARPs product exhibits a positive rate pattern that closely matches that of the definitive product across all training years.
 Figure 16: Comparison of yearly class imbalance in the SHARPs datasets: near-real-time (NRT-SHARPs, with blue colors) versus definitive (DEF-SHARPs, with orange colors).
@@
 Figure 17 presents the average TSS and F1 score gains across eight model configurations.
-Because the NRT-SHARPs dataset is extremely imbalanced during 2023–2024, we compare performance between solar minimum and the combined evolving + solar maximum periods.
+Because the NRT-SHARPs dataset is extremely imbalanced during 2023–2024, we compare performance between solar minimum and the combined evolving plus solar maximum periods.
 During solar minimum, and holding other settings fixed, training on NRT-SHARPs generally improves performance for both logistic and LSTM models.
-In contrast, during the evolving + solar maximum period, NRT-SHARPs decrease performance, particularly in terms of TSS.
+In contrast, during the evolving plus solar maximum period, NRT-SHARPs decrease performance, particularly in terms of TSS.
 The results also demonstrate that models trained on NRT-SHARPs exhibit greater variability than those trained on DEF-SHARPs, likely due to reduced sample quality and a smaller training sample size.
@@
 5 Conclusion
-In this paper, we systematically review and sort the most commonly used data products in solar flare forecasting with machine-learning methods.
+In this paper, we systematically review and compare the most commonly used data products for solar flare forecasting using machine learning methods.
 For the flare-response labels, we cross-evaluate the Science-Quality list from NCEI, the operational SWPC-FTP list, and the SSW catalog.
@@
 In addition, we compare the predictive performance of LSTM-based models and logistic-regression baselines across multiple combinations of flare lists (Science-Quality and SWPC-FTP) and SHARP data (DEF-SHARPs and NRT-SHARPs).
-Our results highlight several important conclusions:
+Our analysis results in the following conclusions.
   1. 1.
@@
   2. 2.
-    NRT-SHARPs introduce substantially greater variability in model skill scores, likely resulting from reduced data quality and a smaller effective training sample.
+    Near-real-time SHARPs, as compared with definitive SHARPs, introduce substantially greater variability in model skill scores, likely resulting from reduced data quality and a smaller effective training sample.
   3. 3.
     The impact of data products is strongly solar-cycle dependent.
-Holding other factors fixed, the operational SWPC-FTP labels improve model performance during solar minimum but degrade performance during solar maximum; NRT-SHARPs exhibit a similar pattern.
+Holding other factors fixed, the operational SWPC-FTP labels improve model performance during solar minimum but degrade performance during solar maximum.
+The NRT-SHARPs and DEF-SHARPs comparison exhibits a similar pattern.
   4. 4.
     Although the more sophisticated LSTM models do not consistently outperform logistic regression, they produce more stable results across forecasting windows and data product combinations, indicating robustness to label and predictor variations.
-By clarifying the strengths and limitations of each product and quantifying their effects on predictive performance, we hope to help researchers in selecting appropriate datasets and to improve the comparability and interpretability of data-driven flare-forecasting studies.
+By clarifying the strengths and limitations of each product and quantifying their effects on predictive performance, we aim to help researchers select appropriate datasets and to improve the comparability and interpretability of data-driven flare-forecasting studies.
 Acknowledgments
@@
 Appendix A Quiet Period Analysis of Active Regions
-In this section, we are trying to clearly identify if there are any distributional differences when an active region has an ongoing flare, versus when it is during a “quiet period”.
-As an initial approach, we are defining the ”quiet period” of an active region to be any period of time that is 24 hours removed from any flare activity (i.e.
-more than 24 hours before the flare start time and after the flare ends).
-We are measuring these distributional differences by looking at the marginal distribution of all the SHARPs depending on if it’s a quiet period or a flaring period (which is defined to be any time that is not a quiet period).
-After running through all the flares from 2010 to 2024 and categorizing the periods of times of all active regions that existed during this interval as quiet or flaring, we analyzed the marginal distriubtion of the SHARPs during this time.
+In this section, we aim to determine whether there are distributional differences between active regions during ongoing flares and during a “quiet period”.
+As an initial approach, we define the ”quiet period” of an active region as any period 24 hours removed from flare activity (i.e., more than 24 hours before the flare start time and after the flare ends).
+We measure these distributional differences by examining the marginal distribution of all SHARPs, depending on whether it’s a quiet period or a flaring period (defined as any time that is not a quiet period).
+After reviewing all flares from 2010 to 2024 and categorizing the periods of all active regions during this interval as quiet or flaring, we analyzed the marginal distribution of SHARPs during this period.
 These are visualized here in Figure 18
@@
 The red curve depicts the exact time interval of a B-flare as described by the GOES flare list, and the purple line expands the interval by 24 hours before and after the start and end times, respectively.
-This plot shows that the distribution of a few SHARPs differs depending on whether it is during a quiet period or when there is activity.
-In particular, the distribution of MEANGAM, MEANSHR, and SHGRT45 suggests that there are indeed differences in the behavior of active regions depending on what time period they are in.
-To better generalize the behavior of active regions, it may be beneficial to incorporate quiet time data of active regions as well.
+This plot shows that the distribution of a few SHARPs differs between quiet periods and periods of activity.
+In particular, the distributions of MEANGAM, MEANSHR, and SHGRT45 suggest that active regions exhibit different behavior across time periods.
+To better generalize the behavior of active regions, it may be beneficial to incorporate quiet-time data from active regions as well.
 Additionally, Figure 19 depicts the class of all flares in the GOES flare list starting from 1996, which encapsulates two solar cycles.
-During the peak of the solar cycle, we can see that a percentage of B-flares is missing.
-This indicates that the flare list is not necessarily capturing all B-flares, as they could be obscured by the stronger flares that occur more frequently during the solar cycle.
-As such, the quiet time imaging data during these solar cycle peaks can also serve as a “representation” of the B-flares during this time.
+At the peak of the solar cycle, a percentage of B-flares is missing.
+This indicates that the flare list may not capture all B-flares, as they could be obscured by stronger flares that occur more frequently during the solar cycle.
+As such, quiet-time imaging data from these solar-cycle peaks can also serve as a representation of the B-flares during this period.
 Figure 19: Marginal Distribution of SHARPs over different periods aggregated across Active Regions.
@@
 If we adjust for this scaling, all the flares from before 2019 will be shifted up the length from the “black” line” to the red line.
-In this scenario, we will have systematically missing low-intensity B-flares in our data, so having the quiet data during this time period can help mitigate any potential negative effects of this.
+In this scenario, we will have systematic missingness of low-intensity B-flares in our data; therefore, having quiet data during this period can help mitigate potential negative effects.
 Appendix B Comparison of Numerical Results